### v1

# ExoJAX Retrievals of VLT/CRIRES Spectra of Luhman 16AB: C/O Ratios and Systematic Uncertainties

Abstract
We present atmospheric retrievals of the benchmark brown dwarf binary Luhman 16AB using high-resolution VLT/CRIRES spectra and the differentiable framework ExoJAX. We derive elemental abundances and temperature–pressure (TTPP) profiles while explicitly testing the robustness of the results against major sources of systematic uncertainty. We first perform retrievals with a power-law TTPP profile and assess the sensitivity of inferred molecular abundances and C/O ratios to different CO line lists (ExoMol, HITEMP with air- and H2-broadening). We then introduce a flexible Gaussian process-based TTPP profile, allowing a non-parametric characterization of the thermal structure and a more conservative treatment of uncertainties. For both components, we infer C/O ratios of about 0.67, slightly above solar, with line list systematics at the 7 percent level emerging as the dominant source of uncertainty, whereas assumptions about TTPP parameterization or photometric variability play a lesser role. The retrieved TTPP profiles and molecular abundances are broadly consistent with atmospheric models and equilibrium chemistry. Our results establish Luhman 16AB as a key anchor for substellar C/O measurements, demonstrate the utility of flexible TTPP modeling in high-resolution retrievals, and highlight the importance of systematic tests — particularly line list uncertainties — for robust comparisons between brown dwarfs and giant exoplanets.
Brown dwarfs (185) — Exoplanet atmospheres (487) — High resolution spectroscopy (2096) — Molecular spectroscopy (2095) — T dwarfs (1679)
facilities: VLT (CRIRES)software: corner (corner), ExoJAX (Kawahara2022; Kawahara2024), FastChem (fastchem; fastchem_cond), JAX (jax2018github), NumPyro (bingham2018pyro; phan2019composable), tinygp (tinygp)
1 Introduction
Detailed characterization of brown dwarfs and exoplanets is becoming possible with high-resolution spectroscopy. In addition to molecular detections with the cross-correlation function, retrieval analyses that directly model spectra to derive molecular abundances and temperature-pressure (TTPP) structures are now commonly applied (e.g., Birkby2018; Snellen2025). Atmospheric properties derived from such retrieval analyses provide valuable insights into the formation pathways of giant planets and brown dwarfs. For example, the relationship between a gas giant’s C/O ratio and its formation environment has long been discussed as an important diagnostic: generally, star-like formation is expected to inherit the elemental ratios of the natal cloud, whereas core accretion within a disk — where solids and gas evolve separately — can alter the initial composition depending on the formation location (e.g., Oberg2011; Espinoza2023). Empirically, C/O ratios measured for wide-orbit gas giants/super-Jupiters from emission spectroscopy have tended to be roughly solar, whereas those inferred for hot Jupiters from transmission spectroscopy, including with JWST, exhibit a wider dispersion, potentially reflecting differences in formation processes (e.g., Walsh2025). While this study focuses on the C/O ratio, elemental abundances such as [C/H] can also provide complementary diagnostics for distinguishing between planetary and brown dwarf formation pathways (e.g., Xuan2024a; Wang2025).
To advance these discussions, brown dwarfs are important comparison targets, providing anchor points in addition to gas giants: while the formation of cold Jupiters remains debated, brown dwarfs likely form in a star-like manner and thus can provide a baseline for C/O. However, the number of available high-resolution measurements remains limited (Figure 1). Since brown dwarfs are often observable at higher S/N than planetary-mass objects, they also offer valuable testbeds for assessing retrieval systematics. Despite rapid progress of atmospheric retrieval, high-resolution retrievals still carry systematic uncertainties from cloud modeling, molecular line lists, TTPP parameterizations, and chemical–equilibrium assumptions, and a community standard has yet to emerge. For example, Picos2024 compared spectral models built on two distinct gradient assumptions of TTPP profiles and showed that the inferred surface gravity can shift appreciably with the chosen TTPP model, while deRegt2024 demonstrated that using alternative $\mathrm{H_{2}O}$ line lists leads to different inferences for the isotopologue ratio $^{12}\mathrm{CO}/^{13}\mathrm{CO}$. Testing their impact with high-quality data is important both for comparative planetology and for understanding these processes themselves.
[…]
The paper is organized as follows. Section 2 describes the data, and Section 3 outlines the modeling framework used for the retrievals. In Section 4, we analyze the line-list dependence within a power-law TTPP framework. Section 5 then introduces a more flexible Gaussian-process TTPP profile, demonstrates its utility, and discusses how assumptions about the TTPP structure propagate into other parameters and their interpretation. Section 6 summarizes our findings, discusses implications, and outlines directions for future work.
Figure 1: Comparison of C/O ratios versus mass for brown dwarfs and directly-imaged super Jupiters based on high-resolution spectra. Literature values and masses are summarized in Table 4; the measurements for Luhman 16A and 16B are from this work. Blue symbols denote objects that are wide-orbit companions, and orange symbols denote isolated objects. For Luhman 16AB, the vertical error bars include the systematic uncertainty from the choice of CO line list quantified in this study.
Figure 2: Retrieval results for the VLT/CRIRES high-resolution spectra of Luhman 16A (left) and Luhman 16B (right), modeled with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP profile. The wavelength coverage is split into four detector segments (chips 1–4). For each chip, the top panel shows the normalized flux data (black points) and the best-fit physical model (blue line). The bottom panel shows the residuals (data minus the best-fit model); orange lines indicate the correlated noise predicted by the GP. Each chip is normalized independently.
2 Data
[…]
3 Retrieval Framework
3.1 Spectrum Model
Table 1: Parameters and Priors for spectral model
Parameter Unit Prior Description
$\log g$ cgs $\mathcal{N}_{[4.66,5.20]}(4.93,0.09)$ (A), $\mathcal{N}_{[4.59,5.13]}(4.86,0.09)$ (B) Surface gravity
RVRV $\mathrm{km}/\mathrm{s}$ $\mathcal{U}(20,30)$ Radial velocity
$\log_{10}\mathrm{VMR_{\mathrm{CO}}}$ - $\mathcal{U}(-6.0,-1.0)$ Volume mixing ratio of $\mathrm{{}^{12}C^{16}O}$
$\log_{10}\mathrm{VMR_{\mathrm{H_{2}O}}}$ - $\mathcal{U}(-6.0,-1.0)$ Volume mixing ratio of $\mathrm{{}^{1}H_{2}^{16}O}$
$\log_{10}\mathrm{VMR_{\mathrm{CH_{4}}}}$ - $\mathcal{U}(-6.0,-1.0)$ Volume mixing ratio of $\mathrm{{}^{12}C^{1}H_{4}}$
$\log_{10}\mathrm{VMR_{\mathrm{HF}}}$ - $\mathcal{U}(-10.0,-5.0)$ Volume mixing ratio of $\mathrm{{}^{1}H^{19}F}$
$\log_{10}P_{\mathrm{c}}$ bar $\mathcal{U}(-2.0,2.0)$ Pressure of cloud center height
$V\sin i$ $\mathrm{~km}/\mathrm{s}$ $\mathcal{U}(10,20)$ (A), $\mathcal{U}(20,30)$ (B) Projected rotational velocity
$q_{1}$ - $\mathcal{U}(0.0,1.0)$ Limb darkening parameter
$q_{2}$ - $\mathcal{U}(0.0,1.0)$ Limb darkening parameter
$\log_{10}a_{\mathrm{GP}}$ - $\mathcal{U}(-4.5,2.0)$ Amplitude of GP
$\log_{10}\tau_{\mathrm{GP}}$ Å $\mathcal{U}(-1.5,0.5)$ Length scale of GP
$A_{1}$ - $\mathcal{U}(1.0,1.2)$ Normalize factor (chip1)
$A_{2}$ - $\mathcal{U}(1.0,1.2)$ Normalize factor (chip2)
$A_{3}$ - $\mathcal{U}(1.0,1.2)$ Normalize factor (chip3)
$A_{4}$ - $\mathcal{U}(1.0,1.2)$ Normalize factor (chip4)
(Power-law TTPP)
$T_{0}$ K $\mathcal{U}(1000,1700)$ Temperature at $1\,\mathrm{bar}$
$\alpha$ - $\mathcal{U}(0.05,0.20)$ Power-law index
(GP TTPP)
$T_{0,\,m}$ K $\mathcal{U}\left(1000,1700\right)$ Temperature at $1\,\mathrm{bar}$ in the GP mean model
$\alpha_{m}$ - $\mathcal{U}(0.05,0.20)$ Power-law index in the GP mean model
$\log_{10}a_{u}$ - $\mathcal{U}\left(-1.0,0.5\right)$ Amplitude of GP TTPP
$\log_{10}\tau_{u}$ - $\mathcal{U}\left(-0.5,1.4\right)$ Length scale for GP TTPP
We modeled the high-resolution spectra of Luhman 16AB with the ExoJAX spectral framework (Kawahara2022; Kawahara2024). ExoJAX solves the radiative transfer by computing, for each pressure interval $\Delta P$ between adjacent levels, the optical–depth increment as the sum of the molecular contribution in a given layer $\sum_{i}\Delta\tau_{\mathrm{mol},i}$, collision–induced absorption (CIA) $\Delta\tau_{\mathrm{CIA}}$, and cloud opacity $\Delta\tau_{\mathrm{c}}$:
$\Delta\tau(\nu)=\sum_{i}\Delta\tau_{\mathrm{mol},i}(\nu)+\Delta\tau_{\mathrm{CIA}}(\nu)+\Delta\tau_{\mathrm{c}},$ (1)
where $\nu$ is the wavenumber. We discretize the atmosphere into 101 pressure levels, logarithmically spaced from $10^{-4}$$10^{2}$ bar. Here, the molecular contribution in a given layer $\Delta\tau_{\mathrm{mol},i}$ is
$\Delta\tau_{\mathrm{mol},i}\left(\nu\right)=\frac{\mathrm{VMR}_{i}}{\mu\,m_{u}\,g}\,\sigma_{i}\left(\nu\right)\,\Delta P,$ (2)
where $\mathrm{VMR}_{i}$ denotes the volume mixing ratio of absorbing species $i$, $m_{u}$ is the atomic mass unit, and $g$ is the surface gravity. The mean molecular weight $\mu$ is calculated self-consistently from the adopted VMRs of all atmospheric components, accounting for both the absorbing species and the bulk gases. The molecular cross section $\sigma_{i}$ is computed from line strengths $S\left(T\right)$ and Voigt profiles $V\left(\nu,\,\beta,\,\gamma_{L}\right)$ as
$\sigma_{i}\left(\nu\right)=\sum_{l}S_{l}\left(T\right)\,V\left(\nu-\hat{\nu},\,\beta_{l}\left(T\right),\,\gamma_{L,\,l}\left(T,\,P\right)\right).$ (3)
Here, $l$ indexes the absorption lines ($l=1,\ldots,N_{l}$) and $\hat{\nu}$ is the line center. ExoJAX evaluates the Voigt profiles only for the lines located within the specified wavelength range, and for each line the Voigt function is computed over that interval. The Gaussian (thermal) width is
$\beta=\sqrt{\frac{k_{\mathrm{B}}T}{M_{i}m_{u}}}\,\frac{\hat{\nu}}{c},$ (4)
[…]
$\Delta\tau_{\mathrm{CIA}}\left(\nu\right)=\left(n_{\mathrm{H_{2}}}^{2}\,\alpha_{\mathrm{H_{2}-H_{2}}}\left(T,\nu\right)+n_{\mathrm{H_{2}}}\,n_{\mathrm{He}}\,\alpha_{\mathrm{H_{2}-He}}\left(T,\nu\right)\right)\,\frac{k_{\mathrm{B}}T}{\mu m_{u}g}\,\frac{\Delta P}{P},$ (5)
where $n_{\mathrm{H_{2}}}$ and $n_{\mathrm{He}}$ are number densities, $\mu$ is the mean molecular weight, and $\alpha_{\mathrm{CIA}}(T,\nu)$ are CIA coefficients taken from the HITRAN database. The cloud optical depth per layer, $\Delta\tau_{\mathrm{c}}$, is modeled as a Gaussian in $\log_{10}P$ with width $w_{\mathrm{c}}$ and central pressure $P_{\mathrm{c}}$:
$\Delta\tau_{\mathrm{c}}=\frac{C}{\sqrt{2\pi w_{\mathrm{c}}^{2}}}\,\exp{\left[-\frac{(\log_{10}P-\log_{10}P_{\mathrm{c}})^{2}}{2w_{\mathrm{c}}^{2}}\right]},$ (6)
where $C$ is the (dimensionless) column optical depth of the cloud. This functional form is motivated by the vertical cloud opacity profiles in the Sonora diamondback models (Morley2024). Given the narrow wavelength range analyzed here, we neglect any wavelength dependence of the cloud optical depth. Because $w_{\mathrm{c}}$ and $P_{\mathrm{c}}$ are strongly degenerate, we fix $w_{\mathrm{c}}=0.3$. To emulate an effectively optically thick gray deck, we fix the column optical depth to a large arbitrary constant, $C=500$.
Using the resulting $\Delta\tau(\nu)$, we solve the radiative transfer with an eight-stream approximation to obtain the local one-dimensional emergent spectrum $S_{\mathrm{raw}}(\nu)$. We then convolve with a rotational broadening kernel computed from the limb-darkening law and the projected rotational velocity $V\sin i$ to obtain the disk-integrated spectrum $F_{\mathrm{raw}}(\nu)$. For limb darkening, we adopt the quadratic law with the $(q_{1},q_{2})$ parametrization as proposed by Kipping2013.
To model the Earth-observed, per-chip normalized spectra, we first apply the Doppler shift due to the brown dwarf radial velocity $RV$. For an intrinsic wavenumber $\nu_{0}$, the observed wavenumber is
$\nu_{\mathrm{obs}}=\nu_{0}\,\frac{c}{c+RV}.$ (7)
Because the observed spectra are normalized per detector, the model is likewise normalized in each wavelength segment. Denoting the model for the four detector segments by $\boldsymbol{F}_{i}$ ($i=1,2,3,4$) and their mean values by $\boldsymbol{\bar{F}}_{i}$, the normalized spectra are
$\boldsymbol{f}_{i}=\frac{\boldsymbol{F}_{i}}{A_{i}\boldsymbol{\bar{F}}_{i}},$ (8)
with $A_{i}$ a per-segment normalization factor. The final model used in the analysis is the concatenation
$\boldsymbol{f}=\left[\boldsymbol{f}_{1},\,\boldsymbol{f}_{2},\,\boldsymbol{f}_{3},\,\boldsymbol{f}_{4}\right].$ (9)
3.2 Inference
ExoJAX is implemented with JAX (jax2018github), a high-performance Python library that supports GPU acceleration and automatic differentiation. Consequently, our spectral model can be evaluated on GPUs and differentiated with respect to arbitrary parameters, enabling gradient-based inference.
Leveraging these features, we infer posterior distributions of the model parameters with Hamiltonian Monte Carlo (HMC; HMC), specifically the No-U-Turn Sampler (NUTS; NUTS) implemented in NumPyro. With prior $p(\boldsymbol{\theta})$, the posterior is given by Bayes’ theorem:
$p(\boldsymbol{\theta}\mid\boldsymbol{d})\propto\mathcal{L}(\boldsymbol{\theta})\,p(\boldsymbol{\theta}).$ (10)
To account for correlated noise, we use a Gaussian-process (GP) likelihood (Kawahara2022),
$\mathcal{L}(\boldsymbol{\theta})=\frac{1}{\sqrt{(2\pi)^{n}\,|\boldsymbol{\Sigma}|}}\,\exp{\left[-\frac{1}{2}\,(\boldsymbol{d}-\boldsymbol{f(\boldsymbol{\theta})})^{\mathrm{T}}\,\boldsymbol{\Sigma}^{-1}\,(\boldsymbol{d}-\boldsymbol{f(\boldsymbol{\theta})})\right]},$ (11)
where $\boldsymbol{d}$ is the $n$-point observed spectrum and $\boldsymbol{f}(\boldsymbol{\theta})$ is the model. The covariance matrix is
$\Sigma_{ij}=K_{ij}+\sigma_{e,\,i}^{2}\,\delta_{ij},$ (12)
with $\sigma_{e,i}$ the $i$-th data uncertainty and $\delta_{ij}$ the Kronecker delta. We adopt an RBF kernel,
$K_{ij}=a_{\mathrm{GP}}\,\exp{\left[-\frac{(\lambda_{i}-\lambda_{j})^{2}}{2\tau_{\mathrm{GP}}^{2}}\right]},$ (13)
where $\lambda_{i},\,\lambda_{j}$ are wavelength of $i$-th and $j$-th data, $a_{\mathrm{GP}}$ is the GP amplitude, and $\tau_{\mathrm{GP}}$ (in $\mathrm{\AA }$) is the correlation length.
The set of inferred parameters and their priors is listed in Table 1. For $\log g$, we imposed Normal priors informed by external constraints. Using the system age of $510\pm 95\,\mathrm{Myr}$ from gagne2023 together with the dynamical masses $M_{A}=33.5\pm 0.3\,\mathrm{M_{J}}$ and $M_{B}=28.6\pm 0.3\,\mathrm{M_{J}}$ from lazorenko2018, we interpolated the brown dwarf evolutionary model of Baraffe2003 to obtain surface gravities. Since the surface gravity is given by $g=GM/R^{2}$ with brown dwarf mass $M$, radius $R$, and the gravitational constant $G$, we propagated the quoted mass uncertainties ($\sim 1\%$) and assumed a 10% fractional uncertainty on the radius to set the prior widths, yielding Normal priors $\log g_{A}=4.93\pm 0.09$ and $\log g_{B}=4.86\pm 0.09$, each truncated at $3\sigma$. Sampling used HMC–NUTS with a dense mass matrix. For numerical stability and mixing, parameters were affinely re-scaled so that the prior support maps to $\mathcal{O}(1)$ in the sampler’s internal coordinates. Sampling/adaptation were performed in the transformed space and results are reported after back-transformation.
4 Retrievals Using a Power-Law TTPP Profile
In this section, we adopt a simple power-law TTPP profile given by
$T\left(P\right)=T_{0}\,\left(\frac{P}{1\,\mathrm{bar}}\right)^{\alpha}.$ (14)
In high-resolution retrievals of brown dwarf and exoplanet atmospheres, line shapes depend on the pressure broadening parameters provided by the molecular line lists, which can in turn impart systematic effects on the inferred molecular abundances and atmospheric structure. To assess the robustness of the results, it is therefore essential to analyze a given target with multiple line lists. Motivated by this, we focus on the CO line list and compare three models based on ExoMol and HITEMP.
4.1 Line Lists
High-resolution retrievals for brown dwarfs commonly use line lists from ExoMol and HITEMP. ExoMol provides theoretically computed line data tailored to brown dwarf and exoplanet atmospheres, assuming an $\mathrm{H_{2}}$$\mathrm{He}$ background. HITEMP, by contrast, is based largely on high-temperature laboratory measurements and typically adopts an Earth background (air) for pressure broadening. More recently, an $\mathrm{H_{2}}$$\mathrm{He}$ broadened variant of HITEMP has become available, obtained by converting the air-broadening parameters; the conversion employs semi-empirical models anchored to validated laboratory measurements and theoretical predictions (tan2022).
4.1.1 ExoMol
According to ExoMol2016, ExoMol supplies transition wavenumbers $\hat{\nu}$, Einstein coefficients $A$, state energies $E$, and statistical weights $\mathsf{g}$, from which the line strength $S(T)$ can be computed as
$S\left(T\right)=\frac{\mathsf{g_{up}}}{Q\left(T\right)}\,\frac{A}{8\pi c\hat{\nu}^{2}}\,e^{-hcE_{\mathrm{low}}/k_{B}T}\,\left(1-e^{-hc\hat{\nu}/k_{B}T}\right),$ (15)
where $E_{\mathrm{low}}$ is the lower-state energy and $\mathsf{g_{up}}$ is the upper-state statistical weight. For pressure broadening in an $\mathrm{H_{2}}$ background, ExoMol provides the reference broadening coefficient $\gamma_{\mathrm{ref}}(J_{\mathrm{low}})$ and temperature exponent $\mathsf{n}(J_{\mathrm{low}})$ as functions of the lower-state rotational quantum number. The pressure-broadened Lorentz width is then
$\gamma^{p}_{L}=\gamma_{\mathrm{ref}}\,\left(\frac{T}{T_{\mathrm{ref}}}\right)^{-\mathsf{n}}\,\left(\frac{P}{P_{\mathrm{ref}}}\right),$ (16)
with $T_{\mathrm{ref}}=296\,\mathrm{K}$ and $P_{\mathrm{ref}}=1\,\mathrm{bar}$.
4.1.2 HITEMP
According to HiTEMP, HITEMP provides $\hat{\nu}$, $S(T_{\mathrm{ref}})$ at $T_{\mathrm{ref}}=296\,\mathrm{K}$, and $E_{\mathrm{low}}$, from which
$S\left(T\right)=S\left(T_{\mathrm{ref}}\right)\,\frac{Q\left(T_{\mathrm{ref}}\right)}{Q\left(T\right)}\,\frac{e^{-hcE_{\mathrm{low}}/k_{\mathrm{B}}T}}{e^{-hcE_{\mathrm{low}}/k_{\mathrm{B}}T_{\mathrm{ref}}}}\,\frac{1-e^{-hc\hat{\nu}/k_{\mathrm{B}}T}}{1-e^{-hc\hat{\nu}/k_{\mathrm{B}}T_{\mathrm{ref}}}}$ (17)
is obtained. HITEMP further provides air-broadened reference widths $\gamma_{\mathrm{ref,\,air}}$ and temperature exponents $\mathsf{n}_{\mathrm{air}}$ for each line, yielding
$\gamma^{p}_{L}=\gamma_{\mathrm{ref,\,air}}\,\left(\frac{T}{T_{\mathrm{ref}}}\right)^{-\mathsf{n}_{\mathrm{air}}}\,\left(\frac{P}{P_{\mathrm{ref}}}\right)$ (18)
with $P_{\mathrm{ref}}=1\,\mathrm{atm}$. For some molecules, $\mathrm{H_{2}}$-broadened parameters $\gamma_{\mathrm{ref,,H_{2}}}$ and $\mathsf{n}_{\mathrm{H_{2}}}$ are available, allowing $\mathrm{H_{2}}$-background pressure broadening to be computed via Eq. (18) (tan2022). In what follows, we refer to air-broadened HITEMP as $\mathrm{HITEMP_{Air}}$ and $\mathrm{H_{2}}$-broadened HITEMP as $\mathrm{HITEMP_{H_{2}}}$. Nominally, air broadening parameters are not appropriate for brown dwarf atmospheres. However, the laboratory experiment by 2025ApJ...984...92H suggests that these parameters are not necessarily incorrect, so we include them in the comparison.
4.1.3 Adopted Line Lists in This Study
Given that brown dwarf atmospheres are dominated by $\mathrm{H_{2}}$, we adopt the ExoMol POKAZATEL for $\mathrm{H_{2}O}$ (POKAZATEL), ExoMol MM for $\mathrm{CH_{4}}$ (MM), and the ExoMol Coxon–Hajig line list for $\mathrm{HF}$ (Coxon-Hajig). To assess the impact of the CO line list — which exhibits prominent absorption in the $K$ band — on the retrieval results, we consider the following three CO line list configurations:
$\mathrm{ExoMol\,\,CO}$
ExoMol CO line list Li2015 (Li2015), assuming an $\mathrm{H_{2}}$ background.
$\mathrm{HITEMP_{Air}\,\,CO}$
HITEMP CO line list (HITEMPCO) with air-broadening parameters.
$\mathrm{HITEMP_{H_{2}}\,\,CO}$
An $\mathrm{H_{2}}$-broadened version of HITEMP CO based on the semi-empirical models constructed on available and validated experimental measurements and theoretical predictions (tan2022).
For all three CO line lists, the same Voigt profile treatment is applied, so that the differences in the retrieval results originate only from the broadening parameters. The selection of molecular species was referenced from the $K$-band high-resolution spectral analysis of Luhman 16AB with IGRINS/Gemini South reported by Ishikawa2025. We also searched for the isotopologue $^{13}\mathrm{CO}$, but it was not detected in the present analysis. This is consistent with Ishikawa2025, who did not detect $^{13}\mathrm{CO}$ in the same wavelength region. Therefore, we did not include $^{13}\mathrm{CO}$ opacity in the spectral model.
4.2 Results
Figure 3: Corner plots for the posterior samples obtained by fitting the spectra with models that use CO line lists from $\mathrm{ExoMol\,\,CO}$ (orange), $\mathrm{HITEMP_{Air}\,\,CO}$ (green), and $\mathrm{HITEMP_{H_{2}}\,\,CO}$ (blue). (Left) Luhman 16A. (Right) Luhman 16B.
Figure 4: Retrieved TTPP profiles and normalized contribution fuctions for Luhman 16A (left) and Luhman 16B (right) under the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration. For each object, the left panel shows the contribution function, and the right panel shows the posterior TTPP profile. The horizontal shaded band marks the inferred cloud deck pressure, $\log_{10}P\left(\tau_{c}=1\right)$. Note that $\tau_{c}$ denotes the cumulative optical depth due only to cloud opacity.
Blue, purple, and red lines are Sonora Diamondback (SD) TTPP models (Morley2024). Black curves in the right panels indicate condensation curves for the principal brown dwarf cloud species identified by Morley2024 (condensation curve formulas from Visscher2010; Wakeford2017).
Table 2: Parameter estimates for Luhman 16A under each modeling configuration.
Parameter Unit PL TTPP $\mathrm{ExoMol\,\,CO}$ PL TTPP $\mathrm{HITEMP_{Air}\,\,CO}$ PL TTPP $\mathrm{HITEMP_{H_{2}}\,\,CO}$ GP TTPP $\mathrm{HITEMP_{H_{2}}\,\,CO}$
$\log g$ cgs $4.93^{+0.07}_{-0.09}$ $5.08^{+0.04}_{-0.06}$ $4.95^{+0.08}_{-0.09}$ $5.06^{+0.08}_{-0.07}$
RVRV $\mathrm{km\,s^{-1}}$ $28.13^{+0.03}_{-0.04}$ $28.19^{+0.05}_{-0.05}$ $28.17^{+0.03}_{-0.04}$ $28.16^{+0.03}_{-0.04}$
$\log_{10}\mathrm{VMR_{CO}}$ $3.20^{+0.06}_{-0.10}$ $3.12^{+0.03}_{-0.04}$ $3.07^{+0.09}_{-0.09}$ $3.04^{+0.07}_{-0.08}$
$\log_{10}\mathrm{VMR_{H_{2}O}}$ $3.47^{+0.08}_{-0.07}$ $49^{+0.03}_{-0.03} 3.3$$6^{+0.09}_{-0.08} 3.3$$31^{+0.08}_{-0.07}$
$\log_{10}\mathrm{VMR_{CH_{4}}}$ $5.17^{+0.10}_{-0.07}$ $5.10^{+0.05}_{-0.05}$ $5.11^{+0.08}_{-0.09}$ $5.04^{+0.08}_{-0.05}$
$\log_{10}\mathrm{VMR_{HF}}$ $7.27^{+0.08}_{-0.09}$ $7.25^{+0.05}_{-0.05}$ $7.17^{+0.07}_{-0.11}$ $7.15^{+0.1}_{-0.07}$
$\mathrm{[C/H]}$ $0.11^{+0.09}_{-0.07}$ $0.19^{+0.03}_{-0.03}$ $0.26^{+0.07}_{-0.09}$ $0.28^{+0.07}_{-0.06}$
$\mathrm{[O/H]}$ $0.06^{+0.08}_{-0.08}$ $0.11^{+0.04}_{-0.03}$ $0.21^{+0.06}_{-0.09}$ $0.23^{+0.07}_{-0.05}$
$\mathrm{[F/H]}$ $3.96^{+0.09}_{-0.08}$ $3.95^{+0.05}_{-0.05}$ $3.85^{+0.09}_{-0.09}$ $3.84^{+0.07}_{-0.07}$
$\mathrm{[(C+O)/H]}$ $0.08^{+0.08}_{-0.08}$ $0.14^{+0.04}_{-0.03}$ $0.23^{+0.06}_{-0.10}$ $0.25^{+0.07}_{-0.05}$
$\mathrm{C/O}$ $0.662^{+0.005}_{-0.006}$ $0.704^{+0.007}_{-0.006}$ $68^{+0.005}_{-0.006} 0.660$$60^{+0.005}_{-0.007}$
$\log_{10}P_{\mathrm{c}}$ bar $1.66^{+0.03}_{-0.06}$ $1.95^{+0.05}_{-0.02}$ $1.52^{+0.02}_{-0.03}$ $1.90^{+0.09}_{-0.05}$
$\log_{10}P\left(\tau_{c}=1\right)$ bar $0.76^{+0.05}_{-0.03}$ $1.05^{+0.05}_{-0.02}$ $0.62^{+0.02}_{-0.03}$ $1.00^{+0.10}_{-0.04}$
$V\sin i$ $\mathrm{km\,s^{-1}}$ $16.21^{+0.37}_{-0.51}$ $15.89^{+0.22}_{-0.29}$ $17.09^{+0.27}_{-0.22}$ $15.62^{+0.19}_{-0.23}$
$q_{1}$ $0.82^{+0.17}_{-0.12}$ $0.73^{+0.26}_{-0.11}$ $0.90^{+0.10}_{-0.05}$ $0.77^{+0.23}_{-0.13}$
$q_{2}$ $0.43^{+0.17}_{-0.15}$ $0.53^{+0.14}_{-0.14}$ $0.29^{+0.09}_{-0.10}$ $0.63^{+0.15}_{-0.12}$
$\log_{10}a_{\mathrm{GP}}$ $3.96^{+0.02}_{-0.02}$ $3.87^{+0.02}_{-0.02}$ $3.95^{+0.01}_{-0.02}$ $4.03^{+0.02}_{-0.02}$
$\log_{10}\tau_{\mathrm{GP}}$ Å $1.27^{+0.01}_{-0.01}$ $1.22^{+0.02}_{-0.01}$ $1.27^{+0.01}_{-0.01}$ $1.30^{+0.01}_{-0.02}$
$A_{1}$ $140^{+0.001}_{-0.001} 1.1$$40^{+0.001}_{-0.001} 1.1$$41^{+0.001}_{-0.001} 1.1$$140^{+0.001}_{-0.001}$
[…]
$A_{3}$ $138^{+0.001}_{-0.001} 1.1$$38^{+0.001}_{-0.001} 1.1$$39^{+0.001}_{-0.001} 1.1$$139^{+0.001}_{-0.001}$
$A_{4}$ $109^{+0.001}_{-0.001} 1.1$$08^{+0.001}_{-0.001} 1.1$$09^{+0.001}_{-0.001} 1.1$$108^{+0.001}_{-0.001}$
(Power-law TTPP)
$T_{0}$ K $1293^{+6}_{-6}$ $1258^{+5}_{-5}$ $1283^{+6}_{-6}$
$\alpha$ $0.109^{+0.003}_{-0.003}$ $0.108^{+0.002}_{-0.003}$ $0.107^{+0.003}_{-0.003}$
(GP TTPP)
$T_{0,\,m}$ K $1285^{+158}_{-143}$
$\alpha_{m}$ $0.110^{+0.027}_{-0.041}$
$\log_{10}a_{u}$ $0.34^{+0.10}_{-0.18}$
$\log_{10}\tau_{u}$ $0.40^{+0.06}_{-0.10}$
(Negative log-likelihood)
Normalized high-resolution spectra 13177-13177 13090-13090 13167-13167 13257-13257
JWST medium-resolution flux density 13 29-29 24-24 1-1
Total 13164-13164 13119-13119 13190-13190 13258-13258
Table 3: Parameter estimates for Luhman 16B under each modeling configuration.
Parameter Unit PL TTPP $\mathrm{ExoMol\,\,CO}$ PL TTPP $\mathrm{HITEMP_{Air}\,\,CO}$ PL TTPP $\mathrm{HITEMP_{H_{2}}\,\,CO}$ GP TTPP $\mathrm{HITEMP_{H_{2}}\,\,CO}$
$\log g$ cgs $4.97^{+0.08}_{-0.06}$ $5.03^{+0.08}_{-0.05}$ $4.96^{+0.07}_{-0.07}$ $5.03^{+0.07}_{-0.05}$
RVRV $\mathrm{km\,s^{-1}}$ $25.66^{+0.07}_{-0.06}$ $25.59^{+0.07}_{-0.07}$ $54^{+0.06}_{-0.08} 25.52$$52^{+0.06}_{-0.06}$
$\log_{10}\mathrm{VMR_{CO}}$ $2.96^{+0.09}_{-0.07}$ $2.97^{+0.09}_{-0.05}$ $2.86^{+0.07}_{-0.09}$ $2.95^{+0.06}_{-0.08}$
$\log_{10}\mathrm{VMR_{H_{2}O}}$ $25^{+0.09}_{-0.05} 3.3$$36^{+0.08}_{-0.05}$ $3.16^{+0.08}_{-0.07}$ $3.21^{+0.06}_{-0.07}$
$\log_{10}\mathrm{VMR_{CH_{4}}}$ $4.65^{+0.08}_{-0.07}$ $4.69^{+0.06}_{-0.06}$ $4.61^{+0.10}_{-0.06}$ $4.61^{+0.07}_{-0.05}$
$\log_{10}\mathrm{VMR_{HF}}$ $7.08^{+0.10}_{-0.07}$ $7.17^{+0.08}_{-0.07}$ $7.00^{+0.10}_{-0.07}$ $7.10^{+0.09}_{-0.07}$
$\mathrm{[C/H]}$ $0.36^{+0.09}_{-0.07}$ $0.35^{+0.06}_{-0.06}$ $0.47^{+0.09}_{-0.08}$ $0.36^{+0.05}_{-0.07}$
$\mathrm{[O/H]}$ $0.30^{+0.09}_{-0.06}$ $0.25^{+0.07}_{-0.05}$ $0.41^{+0.10}_{-0.07}$ $0.31^{+0.06}_{-0.06}$
$\mathrm{[F/H]}$ $3.76^{+0.08}_{-0.09}$ $3.85^{+0.07}_{-0.08}$ $3.69^{+0.07}_{-0.10}$ $3.82^{+0.10}_{-0.07}$
$\mathrm{[(C+O)/H]}$ $0.33^{+0.09}_{-0.06}$ $0.29^{+0.06}_{-0.07}$ $0.43^{+0.09}_{-0.07}$ $0.33^{+0.06}_{-0.07}$
$\mathrm{C/O}$ $0.674^{+0.005}_{-0.008}$ $0.725^{+0.007}_{-0.006}$ $79^{+0.006}_{-0.006} 0.660$$60^{+0.007}_{-0.007}$
$\log_{10}P_{\mathrm{c}}$ bar $1.45^{+0.03}_{-0.03}$ $1.68^{+0.03}_{-0.04}$ $1.35^{+0.03}_{-0.02}$ $1.93^{+0.07}_{-0.03}$
$\log_{10}P\left(\tau_{c}=1\right)$ bar $0.55^{+0.02}_{-0.03}$ $0.78^{+0.03}_{-0.04}$ $0.45^{+0.03}_{-0.02}$ $1.04^{+0.07}_{-0.03}$
$V\sin i$ $\mathrm{km\,s^{-1}}$ $27.65^{+0.28}_{-0.53}$ $27.30^{+0.43}_{-0.37}$ $27.79^{+0.46}_{-0.48}$ $26.93^{+0.56}_{-0.47}$
$q_{1}$ $0.81^{+0.19}_{-0.08}$ $0.81^{+0.18}_{-0.09}$ $0.80^{+0.19}_{-0.09}$ $0.79^{+0.21}_{-0.11}$
$q_{2}$ $0.36^{+0.12}_{-0.11}$ $0.40^{+0.12}_{-0.12}$ $0.34^{+0.11}_{-0.12}$ $0.48^{+0.14}_{-0.14}$
$\log_{10}a_{\mathrm{GP}}$ $3.83^{+0.02}_{-0.02}$ $3.75^{+0.02}_{-0.02}$ $3.83^{+0.02}_{-0.02}$ $3.88^{+0.02}_{-0.02}$
$\log_{10}\tau_{\mathrm{GP}}$ Å $1.32^{+0.01}_{-0.02}$ $1.28^{+0.02}_{-0.02}$ $1.32^{+0.01}_{-0.02}$ $1.35^{+0.01}_{-0.01}$
$A_{1}$ $141^{+0.001}_{-0.001} 1.1$$41^{+0.001}_{-0.001} 1.1$$42^{+0.001}_{-0.001} 1.1$$140^{+0.001}_{-0.001}$
$A_{2}$ $141^{+0.001}_{-0.001} 1.1$$41^{+0.001}_{-0.001} 1.1$$41^{+0.001}_{-0.001} 1.1$$140^{+0.001}_{-0.001}$
$A_{3}$ $141^{+0.001}_{-0.001} 1.1$$40^{+0.001}_{-0.001} 1.1$$42^{+0.001}_{-0.001} 1.1$$142^{+0.001}_{-0.001}$
$A_{4}$ $108^{+0.001}_{-0.001} 1.1$$07^{+0.001}_{-0.001} 1.1$$08^{+0.001}_{-0.001} 1.1$$108^{+0.001}_{-0.001}$
(Power-law TTPP)
$T_{0}$ K $219^{+7}_{-9} 11$$177^{+6}_{-6}$ $1215^{+8}_{-9}$
$\alpha$ $0.129^{+0.003}_{-0.006}$ $0.123^{+0.005}_{-0.004}$ $0.128^{+0.004}_{-0.005}$
(GP TTPP)
$T_{0,\,m}$ K $1165^{+82}_{-143}$
$\alpha_{m}$ $0.104^{+0.023}_{-0.044}$
$\log_{10}a_{u}$ $0.36^{+0.11}_{-0.19}$
$\log_{10}\tau_{u}$ $0.37^{+0.08}_{-0.13}$
(Negative log-likelihood)
Normalized high-resolution spectra 12239-12239 12165-12165 12251-12251 12295-12295
JWST medium-resolution flux density 28-28 14-14 13-13 28-28
Total 12268-12268 12179-12179 12265-12265 12323-12323
Using spectral models based on each CO line list, we performed NUTS sampling and obtained posterior distributions for all parameters. Posterior summaries are listed in Tables 2 and 3.
Figure 2 compares, for both components of Luhman 16, the data and the best-fitting spectral models computed with a power-law TTPP profile, a gray-cloud prescription, and $\mathrm{HITEMP_{H_{2}}\,\,CO}$. Overall, the models reproduce the high-resolution spectra well: the residual panels show no conspicuous absorption features that are clearly misfit. The largest residual structures may reflect limitations of the spectral model (e.g., line list incompleteness or pressure broadening uncertainties), although residual telluric contamination or detector-related systematics cannot be ruled out. Although not shown, the other two CO line lists yield comparably good fits.
To examine posterior structure, Figure 3 overlays selected parameters from the three line list runs in a corner plot. From Eq. (2), the molecular optical depth increment scales as $\Delta\tau_{\mathrm{mol},i}\propto\mathrm{VMR}_{i}/(\mu g)$ (at fixed $T$ and line shape), where the mean molecular weight $\mu$ is computed self-consistently from the adopted VMRs of all species at each sampling step. Since the variation in $\mu$ is negligibly small, its effect on the opacity is minor, and the spectra primarily constrain the ratio $\mathrm{VMR}_{i}/g$. Increases in $g$ can therefore be offset by proportional increases in $\mathrm{VMR}_{i}$, producing the observed positive covariance between $\log g$ and the molecular VMRs. In our spectra the key features of each species are largely separated in wavelength, so the relative line depths directly constrain abundance ratios; accordingly, the retrieved VMRs exhibit positive covariance. These correlations between $\log g$ and the VMRs appear as diagonally elongated contours in the off-diagonal panels of Figure 3. However, this correlation does not translate into a simple one-to-one trend in the one-dimensional summaries (Tables 2 and 3), because the correlation structures in the joint posteriors can also shift due to line-list–dependent systematics and the inferred cloud-deck pressure.
The retrieved TTPP profiles under the power-law + gray-cloud + $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration are shown for each object in Figure 4. Across all line lists, the inferred temperature ranges in the regions of largest contribution are consistent with the Sonora diamondback models (Morley2024) at the literature effective temperatures. Here, $\tau_{c}$ denotes the cumulative optical depth due only to the gray-cloud opacity integrated downward from the top of the atmosphere, and the pressure level at $\tau_{c}=1$ corresponds to the cloud deck in our model. The inferred cloud-deck pressures fall within the altitude ranges predicted by the condensation curves in the Sonora Diamondback models, indicating broad physical consistency with the cloud-base pressures of plausible condensates, although we do not attempt to identify a specific cloud species.
We use the VMR in the retrieval, so that the number density of each species is directly represented as $n_{i}=\mathrm{VMR}_{i}$. From these, we compute the elemental number densities as
$n_{\mathrm{C}}=n_{\mathrm{CO}}+n_{\mathrm{CH_{4}}},$ (19)
$n_{\mathrm{O}}=n_{\mathrm{CO}}+n_{\mathrm{H_{2}O}},$ (20)
$n_{\mathrm{F}}=n_{\mathrm{HF}},$ (21)
$n_{\mathrm{H}}=2n_{\mathrm{H_{2}}}+4n_{\mathrm{CH_{4}}}+2n_{\mathrm{H_{2}O}}+n_{\mathrm{HF}},$ (22)
and the carbon-to-oxygen ratio
$\mathrm{C/O}=\frac{n_{\mathrm{C}}}{n_{\mathrm{O}}}.$ (23)
We also compute $\mathrm{[C/H]}$, $\mathrm{[O/H]}$, $\mathrm{[F/H]}$, and $\mathrm{[(C+O)/H]}$ as
$\mathrm{[X/H]}=\log_{10}\left(\frac{n_{\mathrm{X}}}{n_{\mathrm{H}}}\right)-\log_{10}\left(\mathrm{X/H}\right)_{\odot},\ \mathrm{X}\in\{\mathrm{C,O,F}\},$ (24)
$\mathrm{[(C+O)/H]}=\log_{10}\left(\frac{n_{\mathrm{C}}+n_{\mathrm{O}}}{n_{\mathrm{H}}}\right)-\log_{10}\left(\mathrm{C+O\over H}\right)_{\odot},$ (25)
where the solar ratios $\left(\mathrm{X/H}\right)_{\odot}$ follow Asplund2021. For both Luhman 16A and 16B, the VMR-inferred C/O is $\sim 0.67$ — slightly above solar (Figure 3). We nevertheless find that the inferred C/O shifts by $\sim 7\%$ depending on the adopted line list. Differences in the line shapes alter the inferred cloud-deck pressure and, in turn, the relative VMRs.
To interpret the retrieved abundances, we computed thermochemical-equilibrium VMRs with FastChem including equilibrium condensation (fastchem; fastchem_cond). In these calculations, we adjusted the solar C, O, and F abundances of Asplund2021 by our retrieved $\mathrm{[X/H]}$ values (all other elements fixed to solar). Asplund2021 tabulate $A(\mathrm{X})\equiv\log_{10}(n_{\mathrm{X}}/n_{\mathrm{H}})+12$ with $A(\mathrm{H})=12$; accordingly we adopt
$A(\mathrm{X})=A_{\odot}(\mathrm{X})+\mathrm{[X/H]},\quad\mathrm{X}\in\{\mathrm{C,O,F}\},$ (26)
evaluated separately for Luhman 16A and 16B using their respective $\mathrm{[X/H]}$. Numerically, $A_{\odot}(\mathrm{C})=8.46$, $A_{\odot}(\mathrm{O})=8.69$, and $A_{\odot}(\mathrm{F})=4.40$ (Asplund2021). The equilibrium VMRs computed with these inputs and the retrieved TTPP profiles (Figure 5) agree well for $\mathrm{CO}$, $\mathrm{H_{2}O}$, and $\mathrm{HF}$, indicating compatibility with chemical equilibrium under the inferred thermal structures. By contrast, the retrieved $\mathrm{CH_{4}}$ VMRs for both objects are about an order of magnitude lower than the FastChem predictions; we revisit this discrepancy in Section 5.3.
Figure 5: Retrieved VMRs of $\mathrm{H_{2}O}$ (purple), $\mathrm{CO}$ (brown), $\mathrm{CH_{4}}$ (pink), and $\mathrm{HF}$ (yellow) for Luhman 16A (left) and Luhman 16B (right). Points are plotted at the pressure of the contribution function peak; the vertical extent indicates the pressure range that contributes appreciably for each molecule, with the upper bound truncated at the cloud-deck pressure. For comparison, thermochemical-equilibrium VMRs computed with FastChem, including equilibrium condensation, are shown as 5–95% credible bands.
4.2.1 Comparison with Medium-Resolution Flux Density
Although the retrievals above were performed on normalized spectra, the forward models also provide absolute flux densities. Since observational measurements of the absolute flux are available, checking their consistency with the model is also important for validating the retrieval. We therefore compared the JWST medium-resolution flux density with the model flux densities computed using each CO line list at the chip1 central wavelength $\lambda_{1}=2.30\,\mu\mathrm{m}$ (Figure 6). To enable a direct comparison, we converted the mean emergent flux density per wavenumber from ExoJAX, $\boldsymbol{\bar{F}}_{\mathrm{model}}\,[\mathrm{erg/cm^{2}/s/cm^{-1}}]$, to the observed flux density in JWST units $F_{\mathrm{JWST}}[\mathrm{W/m^{2}/\mu m}]$ via
$F_{\mathrm{JWST}}(\lambda_{1})=\left(\frac{R_{\mathrm{BD}}}{D}\right)^{2}\boldsymbol{\bar{F}}_{\mathrm{model}}\;\frac{10}{\lambda_{1}^{2}},$ (27)
[…]
To assess the agreement across line lists, we compared the total negative log-likelihoods from the combined CRIRES spectra and JWST flux density; the values are shown in the bottom of Table 2 and 3. Among the three CO line lists, $\mathrm{HITEMP_{Air}}$ consistently yielded the poorest fit for both Luhman 16A and B. The remaining two, $\mathrm{ExoMol\,\,CO}$ and $\mathrm{HITEMP_{H_{2}}\,\,CO}$, performed similarly, with $\mathrm{HITEMP_{H_{2}}\,\,CO}$ providing the best agreement for Luhman 16A and nearly matching $\mathrm{ExoMol\,\,CO}$ for Luhman 16B. We therefore focus on the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ results hereafter, noting that our conclusions are insensitive to the choice of line list.
Figure 6: Comparison between the observed JWST medium-resolution flux density (vertical dashed lines) at the central wavelength of chip 1 ($2.30\,\mathrm{\mu m}$) and the model predictions based on different CO line lists (histograms). The model flux densities are converted from ExoJAX emergent fluxes to JWST units using Eq. 27. (Left) Luhman 16A. (Right) Luhman 16B.
4.2.2 Impact of Photometric Variability in Luhman 16B
Luhman 16A shows $2\%$ variability over a 7-hr baseline (Biller2024), and Luhman 16B exhibits $\sim$3% variability over 5 hr (Crossfield2014) and up to $\sim$6% over 7 hr (Biller2024). Such variability, common among L/T transition brown dwarfs including Luhman 16AB, is often attributed to heterogeneous silicate/iron clouds whose evolving rotational modulation produces temporal brightness changes (Radigan2014; Crossfield2014; Chen2024; Saumon_Marley2008). The CRIRES time series used here comprise 14 spectra ($\sim$5 hr) for each component (Crossfield2014) and show comparable normalized flux variability: 2.3% for A and 3.2% for B.
Our baseline retrieval assumes a horizontally homogeneous (i.e., effectively averaged) atmosphere, neglecting cloud patchiness. To assess potential systematics on C/O and cloud‐top pressure from this assumption, we performed independent retrievals for each epoch of Luhman 16B using $\mathrm{HITEMP_{H_{2}}\,\,CO}$. We find no statistically significant differences relative to the stacked‐spectrum results in C/O or cloud deck pressure (Figure 7); quantitatively, the epoch-by-epoch C/O values agree with the stacked value ($\mathrm{HITEMP_{H_{2}}\,\,CO}$) to within at most $1.3\sigma$, and the cloud deck pressures to within at most $2.6\sigma$. Hence, for the present data, the impact of ignoring cloud inhomogeneity appears subdominant to systematics associated with the line lists. This conclusion is consistent with the analysis of the $J$-band VLT/CRIRES+ spectrum by deRegt2025, who found that a patchy-cloud model is only weakly preferred for Luhman 16B and that the retrieval outcomes are largely insensitive to the inclusion of patchiness.
Figure 7: C/O and cloud deck pressure for Luhman 16B from stacked and time-resolved spectra. Left: C/O. Right: cloud deck pressure. In each panel, the region left of the vertical dashed line shows results from stacked spectra using three CO line lists (orange: $\mathrm{ExoMol}$; green: $\mathrm{HITEMP_{Air}}$; blue: $\mathrm{HITEMP_{H_{2}}}$). The region right of the dashed line shows epoch-by-epoch (time-resolved) retrievals using $\mathrm{HITEMP_{H_{2}}\,\,CO}$; from the 14 spectra spanning 0.0–5.0 hr we use five representative epochs. Error bars denote 90% credible region. In the left panel, the dotted horizontal line marks the solar C/O (=0.59=0.59; Asplund2021).
5 Retrievals Using a Flexible Gaussian Process TTPP Profile
Some high-resolution retrieval studies of brown-dwarf and exoplanet atmospheres adopt simple parameterizations of the TTPP profile, such as a power-law form (e.g., Kawahara2022; Kawashima2025; Kasagi2025). Although efficient, such profiles can be overly restrictive. More flexible parameterizations have therefore been proposed, including schemes that parameterize the temperature gradient in discrete pressure layers (Zhang2023) and $P$-splines (e.g., Line2015; deRegt2025). However, these approaches still remain tied to specific functional forms and may fail to capture complex structures.
As a potential future extension that takes advantage of the computational efficiency of HMC/NUTS in high-dimensional settings, [^1] The convergence time of conventional random Metropolis–Hastings scales as $D^{2}$ with $D$ being the number of parameters, whereas HMC scales as $D^{5/4}$ (HMC). There are also studies suggesting that NUTS scales as $D^{1/4}$ when initialized in the concentration region of the canonical Gaussian measure (Bou-Rabee2024). Kawahara2022 proposed a nonparametric TTPP profile model based on Gaussian processes (GP, 2023ARA&A..61..329A). That work, however, did not explore its practical applicability in real retrievals, its sensitivity to other parameters, or its validation through injection–recovery tests. Here we refine the GP-based TTPP profile model into a more practical framework and demonstrate its performance through application to the retrieval of Luhman 16AB data.
5.1 Formulation of the GP TTPP Profile
To flexibly represent temperature–pressure (TTPP) structures while maintaining numerical stability, we define the temperature profile as a GP in a whitened and logistically mapped latent space. This approach ensures physically bounded temperatures, reduces parameter correlations during NUTS sampling, and avoids the need for explicit box priors.
We model the temperature at 21 pressure layers, denoted $\boldsymbol{T}_{\mathrm{GP}}$, as a logistic transformation of a underlying latent variable $\boldsymbol{u}$:
$\boldsymbol{T}_{\mathrm{GP}}=T_{\min}+(T_{\max}-T_{\min})\,\sigma(\boldsymbol{u}),\quad\sigma(\boldsymbol{u})=\frac{1}{1+e^{-\boldsymbol{u}}},$ (28)
where $T_{\min}$ and $T_{\max}$ are the physically allowed temperature limits, set to 210 K and 3500 K, respectively. This transformation guarantees $T_{\mathrm{GP},i}\in[T_{\min},T_{\max}]$ by construction. The latent vector $\boldsymbol{u}$ follows a GP prior with mean function $\boldsymbol{m}_{u}$ and covariance matrix $\boldsymbol{\Sigma}_{u}$:
$\boldsymbol{u}\sim\mathcal{GP}\!\left(\boldsymbol{m}_{u},\,\boldsymbol{\Sigma}_{u}\right).$ (29)
The covariance matrix is defined as
$\Sigma_{u,\,ij}=K_{u,\,ij}+\sigma_{u}^{2}\,\delta_{ij},$ (30)
where we fixed a small jitter term at $\sigma_{u}=10^{-6}$ for numerical stability. This negligible value ensures that the term has no physical impact on the retrieved TTPP profiles. We adopt the Matérn 3/2 kernel,
$K_{u,\,ij}=a_{u}^{2}\left(1+\frac{\sqrt{3}\,|\log_{10}P_{i}-\log_{10}P_{j}|}{\tau_{u}}\right)\exp\!\left[-\frac{\sqrt{3}\,|\log_{10}P_{i}-\log_{10}P_{j}|}{\tau_{u}}\right],$ (31)
with amplitude $a_{u}$ and correlation length $\tau_{u}$ treated as free hyperparameters with uniform priors on $\log_{10}a_{u}$ and $\log_{10}\tau_{u}$. The mean function $\boldsymbol{m}_{u}$ is defined as the logit transform of a baseline temperature profile $\boldsymbol{T}_{\mathrm{mean}}$:
$\boldsymbol{m}_{u}=\log\!\left(\frac{\boldsymbol{T}_{\mathrm{mean}}-T_{\min}}{T_{\max}-T_{\min}}\right)-\log\!\left(1-\frac{\boldsymbol{T}_{\mathrm{mean}}-T_{\min}}{T_{\max}-T_{\min}}\right),$ (32)
where $\boldsymbol{T}_{\mathrm{mean}}$ follow a power-law form
$T_{\mathrm{mean}}(P)=T_{0,\,m}\!\left(\frac{P}{1~\mathrm{bar}}\right)^{\alpha_{m}},$ (33)
with the parameters $T_{0,\,m}$ and $\alpha_{m}$ inferred simultaneously with the GP hyperparameters. For efficient sampling, we adopt the whitened representation:
$\boldsymbol{u}=\boldsymbol{m}_{u}+\mathbf{L}\,\boldsymbol{z},\quad\boldsymbol{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),$ (34)
where $\mathbf{L}$ is the Cholesky factor of the covariance matrix $\boldsymbol{\Sigma}_{u}$, and $\mathbf{I}$ denotes the identity matrix. The whitened variable $\boldsymbol{z}$ is directly sampled in the NUTS inference, which improves convergence and decorrelates the GP components. Finally, the temperature vector $\boldsymbol{T}_{\mathrm{GP}}$ at the 21 layers is interpolated to the 101 layer grid via $\mathrm{IP}(\boldsymbol{T}_{\mathrm{GP}})$. The joint posterior is
$p\!\left(\boldsymbol{\theta},\,\boldsymbol{u},\,T_{0,m},\,\alpha_{m},\,a_{u},\,\tau_{u}\mid\boldsymbol{d}\right)$ $\propto$ $\mathcal{L}\!\left(\boldsymbol{d}\mid\boldsymbol{\theta},\,\mathrm{IP}\!\big(\boldsymbol{T}_{\rm GP}(\boldsymbol{u})\big)\right)\,p\!\left(\boldsymbol{\theta}\right)$ (35)
$\times\;p\!\left(\boldsymbol{u}\mid\boldsymbol{m}_{u}(T_{0,\,m},\alpha_{m}),\,a_{u},\,\tau_{u}\right)\,p\!\left(T_{0,\,m},\,\alpha_{m},\,a_{u},\,\tau_{u}\right),$
which we sample using NUTS.
Relative to the GP formulation of Kawahara2022, our implementation introduces several changes aimed at practical convergence and physical regularization. First, while Kawahara2022 adopted a RBF kernel, we instead use a Matérn 3/2 kernel. The Matérn kernel is less smooth than the RBF kernel, avoiding the overly smooth behavior inherent to the RBF assumption. We also infer the kernel hyperparameters (the amplitude $a_{u}$ and correlation length $\tau_{u}$) jointly with the atmospheric parameters, thereby removing the need to set them manually. In addition, instead of adopting a constant mean function as in Kawahara2022, we use a power-law form characterized by $(T_{0,\,m},\,\alpha_{m})$, whose parameters are inferred jointly with the GP. We find that a constant mean tends to force the retrieved profile to flatten toward the average value in the upper and lower layers, as is also hinted in figure 11 Kawahara2022. The power-law mean model, on the other hand, better reflects the physically expected structure of self-luminous brown dwarfs while allowing the GP to capture deviations from it.
5.2 Demonstration
Figure 8: Retrieval results for the simulation spectra, modeled with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a Gaussian process TTPP profile. The wavelength coverage is split into four detector segments (chips 1–4) like CRIRES. For each chip, the top panel shows the simulated normalized flux data (black points) and the best-fit physical model (red line). The bottom panel shows the residuals (data minus the best-fit model). Each chip is normalized independently.
Figure 9: (Left) The Sonora diamondback (SD) TTPP profile model (black solid line) used to generate the simulation data, and the TTPP profile inferred with the GP TTPP model (red shaded region). (Right) The corresponding corner plot.
We validate the GP TTPP framework using simulated spectra. All physical parameters except those describing the TTPP profile are fixed to the posterior medians from our retrieval of the Luhman 16A spectrum with a power-law TTPP profile and the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list (Table 2). As the input TTPP profile for the simulations, we adopt a Sonora Diamondback model with $T_{\mathrm{eff}}=1300\,\mathrm{K}$, $f_{\mathrm{sed}}=1$, $\log g=5.0$, $\mathrm{M/H}]=1.0$, and solar C/O. The synthetic data $\boldsymbol{d}_{\mathrm{sim}}$ are created by evaluating the spectral model $\boldsymbol{f}(\boldsymbol{\theta})$ at the simulation inputs $\boldsymbol{\theta}_{\mathrm{sim}}$ and adding Gaussian noise with a flux-density uncertainty $\boldsymbol{\sigma}_{\mathrm{sim}}$ matched to the S/N of the Luhman 16A high-resolution data:
$\boldsymbol{d}_{\mathrm{sim}}=\boldsymbol{f}\!\left(\boldsymbol{\theta_{\mathrm{sim}}}\right)+\mathcal{N}\!\left(0,\boldsymbol{\sigma}_{\mathrm{sim}}\right).$ (36)
Because noise is drawn independently at each pixel, the likelihood used in the retrieval is
$\mathcal{L}\!\left(\boldsymbol{\theta}\right)=\prod_{i=1}^{n}\frac{1}{\sqrt{2\pi\sigma_{\mathrm{sim},\,i}^{2}}}\exp\!\left[-\frac{\left(d_{\mathrm{sim},\,i}-f_{i}\!\left(\boldsymbol{\theta}\right)\right)^{2}}{2\sigma_{\mathrm{sim},\,i}^{2}}\right].$ (37)
Priors for the parameters follow Table 1 (shared with Luhman 16A), except for $T_{0,\,m}$, the power-law TTPP parameters, and the GP for the spectral likelihood. We set the prior for $T_{0,\,m}$ to $\mathcal{U}\left(1000,\,2500\right)$ to ensure that the temperature range of the input Sonora Diamondback TTPP profile is fully covered by the allowed parameter space.
The simulation spectra and best fit are shown in Figure 8. The retrieved TTPP profile and posteriors are shown in Figure 9. The GP TTPP model successfully reproduces input physical parameters as well as non–power-law shape in the input TTPP profile at pressures where the contribution function peaks. Conversely, in regions with little or no contribution (e.g., above the photosphere or deep beneath the cloud deck), the temperature uncertainty naturally broadens, reflecting the paucity of information more faithfully than a power-law model. This behavior demonstrates that the GP TTPP approach can capture complex thermal structures where the data are informative, while appropriately inflating uncertainties where they are not.
5.3 Application to Luhman 16AB
Figure 10: Retrieval results for the VLT/CRIRES high-resolution spectra of Luhman 16A (left) and Luhman 16B (right), modeled with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a Gaussian process TTPP profile. The wavelength coverage is split into four detector segments (chips 1–4). For each chip, the top panel shows the normalized flux data (black points) and the best-fit physical model (red line). The middle panel shows the residuals relative to the power-law TTPP model with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list (data minus the power-law TTPP best-fit model), along with the difference between the GP and power-law TTPP best-fit models (red line). The bottom panel shows the residuals relative to the GP TTPP model (data minus the GP TTPP best-fit model); orange lines indicate the correlated noise predicted by the GP. Each chip is normalized independently.
Figure 11: Comparison of posteriors obtained with a power-law TTPP profile (blue) and with a GP TTPP profile (red), both using the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list. (Left) Luhman 16A. (Right) Luhman 16B.
Figure 12: Retrieved TTPP profiles and contribution functions: GP (red shaded) versus power-law (blue) under the gray-cloud + $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration. (Left) Luhman 16A. (Right) Luhman 16B.
Figure 13: Comparison between the retrieved volume mixing ratios and FastChem equilibrium predictions (solar abundances from Asplund2021). (Left) Luhman 16A. (Right) Luhman 16B.
We applied the GP TTPP retrieval to the CRIRES data of Luhman 16AB. Best-fit model spectra for both components are shown in Figure 10. Corner plots for the selected parameters are shown in Figure 11; the retrieved TTPP profiles are shown in Figure 12 along with the power-law profiles from Section 4; the retrieved parameter values are summarized in the rightmost columns of Tables 2 and 3.
Figure 12 shows that the broad trends of the inferred TTPP profiles are consistent with those obtained under the power-law parameterization. Consistently, the molecular abundances and C/O ratios agree with the power-law results within uncertainties, indicating that these bulk parameters are robust to the choice of TTPP profile. As in the simulation test, the temperature uncertainties increase at pressures where the spectral contribution is small, yielding a more conservative characterization than the power-law model.
At the same time, the GP TTPP profiles infer departures from a power-law in the detailed structure. In both components, we find evidence for localized inversions near $\log_{10}P\sim 0.5$ and $\log_{10}P\sim-1$. The feature around $\log_{10}P\sim 0.5$ roughly coincides with the cloud-top pressure inferred in the power-law analysis; an approximately isothermal layer can mimic the behavior of a gray cloud, underscoring that cloud-top inferences are sensitive to the flexibility allowed in the TTPP profile. The higher-altitude feature near $\log_{10}P\sim-1$ corresponds to subtle differences in the depths of line cores arising from cooler layers. Whether this structure is physical remains uncertain; possibilities include residual line-list/systematic effects or additional weak condensate/haze opacity at those pressures. Indeed, Sonora models suggest some condensates may be expected in this regime depending on composition and thermal structure. Regardless, the information content from these levels is limited, so any model deficiency there has a negligible impact on global parameters such as C/O for the present data.
Figure 13 shows the comparison with FastChem, analogous to the power-law case. While the overall trends remain similar, the model-predicted ranges broaden owing to the propagated TTPP uncertainty. The effect is most pronounced for $\mathrm{CH_{4}}$ in deeper layers, for which the apparent discrepancy seen under the power-law assumption is no longer compelling once the increased structural uncertainty is accounted for. This highlights that accurately characterizing the uncertainty in regions with modest contribution is essential when assessing possible departures from chemical equilibrium. In this regard, the GP TTPP profile provides a useful and conservative framework.
6 Summary and Discussion
Using ExoJAX, we performed high-resolution retrievals of Luhman 16AB under a gray-cloud assumption and inferred elemental abundance ratios. Taking the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP parameterization as a baseline, we then (i) repeated the analysis with alternative CO line lists and (ii) introduced a more flexible TTPP description based on GPs to assess the robustness of the inferred parameters.
6.1 Comparison with Previous Work
In addition to the present work, several other high-resolution retrieval studies of Luhman 16AB have recently been reported. Since the spectral wavelength ranges and model setups differ across these works, we focus here on comparing the retrieved physical parameters, namely, molecular abundances and cloud parameters.
Abundance
The molecules included in our model, H2O, CH4, CO, and HF, have all been detected in the $H$- and $K$-band Gemini South/IGRINS high-resolution ($R\sim 28,000$) spectra of Luhman 16AB presented by Ishikawa2025. Their study compared the spectra against forward models generated with PICASO, which incorporates radiative–convective equilibrium, Sonora Bobcat TTPP profiles, VIRGA cloud models, and nonequilibrium chemistry parameterized by a vertical eddy diffusion coefficient $K_{zz}$. The retrieved volume mixing ratios (VMRs) of the common molecules (their Figure 8) are broadly consistent with ours, but tend to be smaller, likely because they fixed the metallicity to solar whereas we inferred slightly super-solar values. Both H2O and HF have also been detected in the CRIRES+/VLT $J$-band high-resolution retrieval by deRegt2025. Their modeling adopted a TTPP parameterization based on the temperature gradient at five knots (as in Zhang2023), and clouds were treated as gray opacity decreasing with a power-law of pressure from the cloud base upward. Despite these differences, the retrieved molecular abundances are consistent with ours. Neither study provides explicit estimates of the C/O ratio: Ishikawa2025 fixed the metallicity and C/O to solar, and in the spectral range analyzed by deRegt2025, no carbon-bearing molecules were detected.
Cloud
Ishikawa2025 also detected H2 lines. By comparing the inferred H2 column density with the interstellar gas-to-dust ratio, they suggested that dust clouds may reside deeper than the H2 absorption layer, located at $\sim 3\,\mathrm{bar}$ (or $\log_{10}P\sim 0.5$) in both components A and B. Their self-consistent models likewise imply that clouds are present at similar or deeper pressures relative to the H2 line-forming region. In the power-law cloud model of deRegt2025, the cloud opacity is dominated by contributions around $\log_{10}P\sim 0.5$–1, and this estimate is robust to assumptions about patchiness. In our analysis as well, the inferred cloud top pressures consistently fall within $\log_{10}P\sim 0.5$–1, broadly agreeing with these results. On the other hand, the quantitative estimates depend on assumptions about the TTPP profile and line lists, which in turn propagate into the inferred C/O ratio as discussed in the next subsection.
6.2 $\mathrm{C/O}$ Ratios
Figure 14 compares $\mathrm{C/O}$ inferred with each CO line list configuration and with the GP TTPP model (using $\mathrm{HITEMP_{H_{2}}\,\,CO}$) to literature values for other brown dwarfs and super Jupiters. For both A and B, the choice of CO line list introduces a systematic offset of $\sim$7% in $\mathrm{C/O}$. By contrast, the assumed TTPP profile (power-law versus GP) does not yield a significant difference in $\mathrm{C/O}$. An epoch-by-epoch analysis of the Luhman 16B time series likewise reveals no statistically significant impact from cloud-cover variability on the retrieved $\mathrm{C/O}$ or molecular VMRs (Section 4.2.2). We therefore conclude that, in our framework, the principal systematic on $\mathrm{C/O}$ arises from the CO line list, whereas the TTPP shape and photometric variability are not dominant. In all cases, the C/O values for components A and B agree within $\approx 2\sigma$, consistent with the expectation that isolated brown dwarfs (and their binaries) form in a star-like manner.
At face value, Figure 14 hints that brown dwarfs may exhibit higher $\mathrm{C/O}$ than super Jupiters. Although this is a collection of inhomogeneous inferences from the literature, our measurements for Luhman 16AB, together with the associated systematic-error budget, reinforce this possibility. While differences in formation pathways (i.e., gravitational instability vs. core accretion) remain a possible explanation, the contrast could also reflect differences in the efficiency of oxygen sequestration by silicate (and other) clouds, as emphasized in previous low-resolution studies (e.g., Line2015; Line2017): more efficient rainout in field brown dwarfs would raise the gas-phase C/O, whereas weaker sequestration in super Jupiters would leave C/O closer to the stellar value. A decisive test will require a larger, uniformly analyzed sample of both planets and brown dwarfs, so that inter-study systematics are minimized and C/O can be compared on equal footing. Building on such a uniform baseline, joint analyses of high-resolution spectra (to constrain TTPP) and JWST coverage (to probe silicate or other cloud composition and vertical structure) can further mitigate cloud–TTPP–chemistry degeneracies, thereby providing a more direct test of the efficiency of oxygen sequestration by rainout. We return to the prospects and implementation of such joint retrievals in Section 6.4.
Figure 14: Comparison of $\mathrm{C/O}$ for Luhman 16AB derived in this work with $\mathrm{C/O}$ values reported for other brown dwarfs and directly-imaged super Jupiters from high-resolution spectroscopic analyses (Table 4). Abbreviations: PLTP = power-law TTPP profile; GPTP = Gaussian-process TTPP profile.
6.3 Usefulness and Caveats of the GP TTPP Model
For both simulations and real data, the GP TTPP model captures structures that are more complex than a simple power-law and yields a more faithful assessment of uncertainties. In the present application, these differences have only a modest impact on bulk parameters such as $\mathrm{C/O}$. However, as discussed in Section 5.3, comparisons between retrieved molecular abundances and chemical-equilibrium predictions are sensitive to how accurately TTPP uncertainties are estimated; in this context, the GP framework is important because it enables predictive uncertainties to faithfully reflect the spectra’s information content, remaining conservative where the data are weakly informative. The GP approach is also sensitive to residual model–data mismatches and thus shows promise for diagnosing deficiencies in current models, potentially pointing to additional physics.
On the other hand, as indicated in Section 5.3, degeneracies between the TTPP shape and cloud opacity can also allow nonphysical temperature structures if left unconstrained. Mitigations include imposing weak physical regularization (e.g., monotonicity or stability constraints) on flexible TTPP profiles and/or extending the analysis to broader wavelength coverage where cloud signatures are stronger. The latter strategy can help balance flexibility with physical plausibility; see also Section 6.4.
Figure 15: Comparison between the medium-resolution spectra predicted by extrapolating our high-resolution retrievals and the JWST observations. Shaded bands show the 90% credible intervals of the predicted spectra computed with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP profile (blue: Luhman 16A; orange: Luhman 16B). Points show the corresponding JWST medium-resolution spectra from Biller2024; Chen2025 (blue: A; orange: B).
6.4 Cloud Degeneracies and Joint Analysis with JWST
Throughout this work, we find that the inferred cloud-top pressure is sensitive to assumptions about the TTPP profile and the adopted line list. These systematics propagate into the retrieved VMRs and, via absolute flux, into radius inferences. This underscores the value of combining broad wavelength information with high spectral resolution.
[…]
Building upon these insights, future work will incorporate more detailed cloud modeling that accounts for realistic particle size distributions, compositions, and vertical structures. While the present analysis adopted a simple gray cloud prescription (Eq. 6) to focus on the line list and TTPP effects, these developments — particularly when combined with joint retrievals using JWST and high-resolution ground-based spectra — will enable a more comprehensive understanding of cloud formation and evolution in substellar atmospheres.
We gratefully acknowledge Ian J. M. Crossfield for providing the reduced VLT/CRIRES spectra of Luhman 16AB. Based on observations collected at the European Southern Observatory under ESO programme(s) 291.C-5006(A). We also thank the anonymous reviewer for providing comments that helped improve the manuscript. This study was supported by JSPS KAKENHI Grant Numbers 24KJ1555 (H.Y.), 21H04998 (K.M., Y.K., and H.K.), 22H05150 (Y.K.), 25K17429 (Y.K.), 23H01224 (Y.K. and H.K.), and 23H00133 (H.K.). We used ChatGPT (OpenAI, GPT-5, Sept 2025) for language editing. No part of the scientific analysis was performed by ChatGPT. The authors are responsible for the content and interpretation. Table 4 provides the literature values, together with the corresponding references, for the objects shown in Figure 1.
Table 4: Literature C/O Ratios and Masses for Brown Dwarfs and Super Jupiters
Object ClassaaClass: FBD = Field Brown Dwarf; YBD = Young Brown Dwarf ($\lesssim\text{few}\times 10\,\mathrm{Myr}$); SJ = Super Jupiter Comp.bbCompanion flag: Y = bound to a host star (long-period); N = isolated. C/O $M$ $\mathrm{M_{J}}$] $T_{\mathrm{eff}}$ [K] Ref.
DENIS J0255 FBD N $0.681\pm 0.005$ $45.7\pm 4.4$ $1554^{+263}_{-53}$ (1)
HD 4747 B FBD Y $0.66\pm 0.04$ $67.2\pm 1.8$ $1554^{+128}_{-218}$ (2)
HIP 55507 B FBD Y $0.67\pm 0.04$ $88.0^{+3.4}_{-3.2}$ $2350\pm 50$ (3)
HD 33632 Ab FBD Y $0.58\pm 0.14$ $46\pm 8$ $1628^{+109}_{-97}$ (4), (5)
Gl 229 B FBD Y $0.77\pm 0.01$ $71.47^{+0.57}_{-0.61}$ccGl 229 B has been reported to be a binary, based on high–angular-resolution observations with GRAVITY/VLTI together with radial-velocity monitoring with CRIRES+/VLT (Xuan2024b). $850\pm 50$ (6), (7)
HR 7672 B FBD Y $0.52\pm 0.02$ $72.84^{+0.71}_{-0.72}$ $1806\pm 77$ (8)
2MASS J0835 FBD N $0.65\pm 0.02$ $62.47\pm 15.86$ $1754\pm 112$ (9), (10)
2MASS J0501 FBD N $0.68\pm 0.04$ $21.45\pm 13.71$ $1720\pm 55$ (9), (11)
2MASS J0500 FBD N $0.71\pm 0.01$ $63.68\pm 14.44$ $1793\pm 72$ (9), (12)
HD 984 B YBD Y $0.50\pm 0.01$ $61\pm 4$ $2730^{+120}_{-180}$ (13), (14), (15)
TWA 28 YBD N $0.61\pm 0.02$ $20.9\pm 5.0$ $2382\pm 42$ (16), (17), (18)
2MASS J0856 YBD N $0.58\pm 0.01$ $14.4\pm 1.4$ $2380\pm 32$ (16), (18), (19)
2MASS J0355 YBD N $0.56\pm 0.02$ $19^{+7}_{-5}$ $1527\pm 14$ (20), (21)
HIP 99770 b SJ Y $0.55\pm 0.05$ $16.1^{+5.4}_{-5.0}$ $1400^{+200}_{-150}$ (22), (23)
$\beta$ Pic b SJ Y $0.41\pm 0.05$ $11.9\pm 3.0$ $1724\pm 15$ (24), (25), (26)
GQ Lup b SJ Y $0.50\pm 0.01$ $25^{+11}_{-15}$ $2650\pm 100$ (27), (28)
YSES 1 b SJ Y $0.57\pm 0.01$ $21.8\pm 3.0$ $1893\pm 10$ (29), (30)
YSES 1 c SJ Y $0.36\pm 0.15$ $7.2\pm 0.7$ $958^{+163}_{-76}$ (29), (30)
$\kappa$ And b SJ Y $0.58\pm 0.05$ $22\pm 9$ $1680^{+60}_{-50}$ (31)
GSC 6214 b SJ Y $0.70\pm 0.07$ $21\pm 6$ $1860^{+170}_{-110}$ (31)
ROXs 42 B SJ Y $0.48\pm 0.08$ $13\pm 5$ $2270^{+170}_{-140}$ (31)
DH Tau b SJ Y $0.54\pm 0.06$ $12\pm 4$ $2250^{+120}_{-100}$ (31)
ROXs 12 b SJ Y $0.54\pm 0.05$ $19\pm 5$ $2500\pm 140$ (31)
HIP 79098 b SJ Y $0.54\pm 0.03$ $28\pm 13$ $2360^{+70}_{-80}$ (31)
2MASS J01225 SJ Y $0.37\pm 0.08$ $25\pm 12$ $1710^{+170}_{-160}$ (31)
AB Pic b SJ Y $0.59\pm 0.01$ $30^{+20}_{-18}$ $2000^{+100}_{-300}$ (32), (33)
Note. — References—(1) deRegt2024; (2) Xuan2022; (3) Xuan2024c; (4) Hsu2024; (5) Currie2020; (6) Kawashima2025; (7) Calamari2022; (8) Wang2022; (9) Mulder2025; (10) Schlawin2017; (11) Zapatero2014; (12) Filippazzo2015; (13) Costes2024; (14) Franson2022; (15) Johnson-Groh2017; (16) Picos2024; (17) Scholz2005; (18) Cooper2024; (19) Morales2021; (20) Zhang2021; (21) Vos2022; (22) Zhang2024b; (23) Currie2023; (24) Landman2024; (25) Lacour2021; (26) Chilcote2017; (27) Picos2025; (28) Schwarz2016; (29) Zhang2024a; (30) Wood2023; (31) Xuan2024a; (32) Gandhi2025; (33) Bonnefoy2010

### v2

@@
 Brown dwarfs (185) — Exoplanet atmospheres (487) — High resolution spectroscopy (2096) — Molecular spectroscopy (2095) — T dwarfs (1679)
-facilities: VLT (CRIRES)software: corner (corner), ExoJAX (Kawahara2022; Kawahara2024), FastChem (fastchem; fastchem_cond), JAX (jax2018github), NumPyro (bingham2018pyro; phan2019composable), tinygp (tinygp)
-1 Introduction
+facilities: VLT (CRIRES)software: corner (Foreman-Mackey, 2016), ExoJAX (Kawahara et al., 2022, 2025), FastChem (Stock et al., 2018; Kitzmann et al., 2024), JAX (Bradbury et al., 2018), NumPyro (Bingham et al., 2018; Phan et al., 2019), tinygp (Foreman-Mackey et al., 2024)
+I Introduction
 Detailed characterization of brown dwarfs and exoplanets is becoming possible with high-resolution spectroscopy.
-In addition to molecular detections with the cross-correlation function, retrieval analyses that directly model spectra to derive molecular abundances and temperature-pressure (TTPP) structures are now commonly applied (e.g., Birkby2018; Snellen2025).
+In addition to molecular detections with the cross-correlation function, retrieval analyses that directly model spectra to derive molecular abundances and temperature-pressure (TTPP) structures are now commonly applied (e.g., Birkby, 2018; Snellen, 2025).
 Atmospheric properties derived from such retrieval analyses provide valuable insights into the formation pathways of giant planets and brown dwarfs.
-For example, the relationship between a gas giant’s C/O ratio and its formation environment has long been discussed as an important diagnostic: generally, star-like formation is expected to inherit the elemental ratios of the natal cloud, whereas core accretion within a disk — where solids and gas evolve separately — can alter the initial composition depending on the formation location (e.g., Oberg2011; Espinoza2023).
-Empirically, C/O ratios measured for wide-orbit gas giants/super-Jupiters from emission spectroscopy have tended to be roughly solar, whereas those inferred for hot Jupiters from transmission spectroscopy, including with JWST, exhibit a wider dispersion, potentially reflecting differences in formation processes (e.g., Walsh2025).
-While this study focuses on the C/O ratio, elemental abundances such as [C/H] can also provide complementary diagnostics for distinguishing between planetary and brown dwarf formation pathways (e.g., Xuan2024a; Wang2025).
+For example, the relationship between a gas giant’s C/O ratio and its formation environment has long been discussed as an important diagnostic: generally, star-like formation is expected to inherit the elemental ratios of the natal cloud, whereas core accretion within a disk — where solids and gas evolve separately — can alter the initial composition depending on the formation location (e.g., Öberg et al., 2011).
+Empirically, C/O ratios measured for wide-orbit gas giants/super-Jupiters from emission spectroscopy have tended to be roughly solar, whereas those inferred for hot Jupiters from transmission spectroscopy, including with JWST, exhibit a wider dispersion, potentially reflecting differences in formation processes (e.g., Walsh, 2025).
+While this study focuses on the C/O ratio, elemental abundances such as [C/H] can also provide complementary diagnostics for distinguishing between planetary and brown dwarf formation pathways (e.g., Xuan et al., 2024a; Wang, 2025).
 To advance these discussions, brown dwarfs are important comparison targets, providing anchor points in addition to gas giants: while the formation of cold Jupiters remains debated, brown dwarfs likely form in a star-like manner and thus can provide a baseline for C/O.
@@
 Despite rapid progress of atmospheric retrieval, high-resolution retrievals still carry systematic uncertainties from cloud modeling, molecular line lists, TTPP parameterizations, and chemical–equilibrium assumptions, and a community standard has yet to emerge.
-For example, Picos2024 compared spectral models built on two distinct gradient assumptions of TTPP profiles and showed that the inferred surface gravity can shift appreciably with the chosen TTPP model, while deRegt2024 demonstrated that using alternative $\mathrm{H_{2}O}$ line lists leads to different inferences for the isotopologue ratio $^{12}\mathrm{CO}/^{13}\mathrm{CO}$.
+For example, González Picos et al.
+(2024) compared spectral models built on two distinct gradient assumptions of TTPP profiles and showed that the inferred surface gravity can shift appreciably with the chosen TTPP model, while de Regt et al.
+(2024) demonstrated that using alternative $\mathrm{H_{2}O}$ line lists leads to different inferences for the isotopologue ratio $^{12}\mathrm{CO}/^{13}\mathrm{CO}$.
 Testing their impact with high-quality data is important both for comparative planetology and for understanding these processes themselves.
@@
 The paper is organized as follows.
-Section 2 describes the data, and Section 3 outlines the modeling framework used for the retrievals.
-In Section 4, we analyze the line-list dependence within a power-law TTPP framework.
-Section 5 then introduces a more flexible Gaussian-process TTPP profile, demonstrates its utility, and discusses how assumptions about the TTPP structure propagate into other parameters and their interpretation.
-Section 6 summarizes our findings, discusses implications, and outlines directions for future work.
+Section II describes the data, and Section III outlines the modeling framework used for the retrievals.
+In Section IV, we analyze the line-list dependence within a power-law TTPP framework.
+Section V then introduces a more flexible Gaussian-process TTPP profile, demonstrates its utility, and discusses how assumptions about the TTPP structure propagate into other parameters and their interpretation.
+Section VI summarizes our findings, discusses implications, and outlines directions for future work.
 Figure 1: Comparison of C/O ratios versus mass for brown dwarfs and directly-imaged super Jupiters based on high-resolution spectra.
@@
 Each chip is normalized independently.
-2 Data
+II Data
 […]
-3 Retrieval Framework
-3.1 Spectrum Model
+III Retrieval Framework
+III.1 Spectrum Model
 Table 1: Parameters and Priors for spectral model
@@
 $\log_{10}\tau_{u}$ - $\mathcal{U}\left(-0.5,1.4\right)$ Length scale for GP TTPP
-We modeled the high-resolution spectra of Luhman 16AB with the ExoJAX spectral framework (Kawahara2022; Kawahara2024).
+We modeled the high-resolution spectra of Luhman 16AB with the ExoJAX spectral framework (Kawahara et al., 2022, 2025).
 ExoJAX solves the radiative transfer by computing, for each pressure interval $\Delta P$ between adjacent levels, the optical–depth increment as the sum of the molecular contribution in a given layer $\sum_{i}\Delta\tau_{\mathrm{mol},i}$, collision–induced absorption (CIA) $\Delta\tau_{\mathrm{CIA}}$, and cloud opacity $\Delta\tau_{\mathrm{c}}$:
@@
 where $C$ is the (dimensionless) column optical depth of the cloud.
-This functional form is motivated by the vertical cloud opacity profiles in the Sonora diamondback models (Morley2024).
+This functional form is motivated by the vertical cloud opacity profiles in the Sonora diamondback models (Morley et al., 2024).
 Given the narrow wavelength range analyzed here, we neglect any wavelength dependence of the cloud optical depth.
@@
 We then convolve with a rotational broadening kernel computed from the limb-darkening law and the projected rotational velocity $V\sin i$ to obtain the disk-integrated spectrum $F_{\mathrm{raw}}(\nu)$.
-For limb darkening, we adopt the quadratic law with the $(q_{1},q_{2})$ parametrization as proposed by Kipping2013.
+For limb darkening, we adopt the quadratic law with the $(q_{1},q_{2})$ parametrization as proposed by Kipping (2013).
 To model the Earth-observed, per-chip normalized spectra, we first apply the Doppler shift due to the brown dwarf radial velocity $RV$.
@@
 $\boldsymbol{f}=\left[\boldsymbol{f}_{1},\,\boldsymbol{f}_{2},\,\boldsymbol{f}_{3},\,\boldsymbol{f}_{4}\right].$ (9)
-3.2 Inference
-ExoJAX is implemented with JAX (jax2018github), a high-performance Python library that supports GPU acceleration and automatic differentiation.
+III.2 Inference
+ExoJAX is implemented with JAX (Bradbury et al., 2018), a high-performance Python library that supports GPU acceleration and automatic differentiation.
 Consequently, our spectral model can be evaluated on GPUs and differentiated with respect to arbitrary parameters, enabling gradient-based inference.
-Leveraging these features, we infer posterior distributions of the model parameters with Hamiltonian Monte Carlo (HMC; HMC), specifically the No-U-Turn Sampler (NUTS; NUTS) implemented in NumPyro.
+Leveraging these features, we infer posterior distributions of the model parameters with Hamiltonian Monte Carlo (HMC; Neal, 2011), specifically the No-U-Turn Sampler (NUTS; Hoffman and Gelman, 2014) implemented in NumPyro.
 With prior $p(\boldsymbol{\theta})$, the posterior is given by Bayes’ theorem:
 $p(\boldsymbol{\theta}\mid\boldsymbol{d})\propto\mathcal{L}(\boldsymbol{\theta})\,p(\boldsymbol{\theta}).$ (10)
-To account for correlated noise, we use a Gaussian-process (GP) likelihood (Kawahara2022),
+To account for correlated noise, we use a Gaussian-process (GP) likelihood (Kawahara et al., 2022),
 $\mathcal{L}(\boldsymbol{\theta})=\frac{1}{\sqrt{(2\pi)^{n}\,|\boldsymbol{\Sigma}|}}\,\exp{\left[-\frac{1}{2}\,(\boldsymbol{d}-\boldsymbol{f(\boldsymbol{\theta})})^{\mathrm{T}}\,\boldsymbol{\Sigma}^{-1}\,(\boldsymbol{d}-\boldsymbol{f(\boldsymbol{\theta})})\right]},$ (11)
@@
 For $\log g$, we imposed Normal priors informed by external constraints.
-Using the system age of $510\pm 95\,\mathrm{Myr}$ from gagne2023 together with the dynamical masses $M_{A}=33.5\pm 0.3\,\mathrm{M_{J}}$ and $M_{B}=28.6\pm 0.3\,\mathrm{M_{J}}$ from lazorenko2018, we interpolated the brown dwarf evolutionary model of Baraffe2003 to obtain surface gravities.
+Using the system age of $510\pm 95\,\mathrm{Myr}$ from Gagné et al.
+(2023) together with the dynamical masses $M_{A}=33.5\pm 0.3\,\mathrm{M_{J}}$ and $M_{B}=28.6\pm 0.3\,\mathrm{M_{J}}$ from Lazorenko and Sahlmann (2018), we interpolated the brown dwarf evolutionary model of Baraffe et al.
+(2003) to obtain surface gravities.
 Since the surface gravity is given by $g=GM/R^{2}$ with brown dwarf mass $M$, radius $R$, and the gravitational constant $G$, we propagated the quoted mass uncertainties ($\sim 1\%$) and assumed a 10% fractional uncertainty on the radius to set the prior widths, yielding Normal priors $\log g_{A}=4.93\pm 0.09$ and $\log g_{B}=4.86\pm 0.09$, each truncated at $3\sigma$.
@@
 Sampling/adaptation were performed in the transformed space and results are reported after back-transformation.
-4 Retrievals Using a Power-Law TTPP Profile
+IV Retrievals Using a Power-Law TTPP Profile
 In this section, we adopt a simple power-law TTPP profile given by
@@
 Motivated by this, we focus on the CO line list and compare three models based on ExoMol and HITEMP.
-4.1 Line Lists
+IV.1 Line Lists
 High-resolution retrievals for brown dwarfs commonly use line lists from ExoMol and HITEMP.
@@
 HITEMP, by contrast, is based largely on high-temperature laboratory measurements and typically adopts an Earth background (air) for pressure broadening.
-More recently, an $\mathrm{H_{2}}$$\mathrm{He}$ broadened variant of HITEMP has become available, obtained by converting the air-broadening parameters; the conversion employs semi-empirical models anchored to validated laboratory measurements and theoretical predictions (tan2022).
-4.1.1 ExoMol
-According to ExoMol2016, ExoMol supplies transition wavenumbers $\hat{\nu}$, Einstein coefficients $A$, state energies $E$, and statistical weights $\mathsf{g}$, from which the line strength $S(T)$ can be computed as
+More recently, an $\mathrm{H_{2}}$$\mathrm{He}$ broadened variant of HITEMP has become available, obtained by converting the air-broadening parameters; the conversion employs semi-empirical models anchored to validated laboratory measurements and theoretical predictions (Tan et al., 2022).
+IV.1.1 ExoMol
+According to Tennyson et al.
+(2016), ExoMol supplies transition wavenumbers $\hat{\nu}$, Einstein coefficients $A$, state energies $E$, and statistical weights $\mathsf{g}$, from which the line strength $S(T)$ can be computed as
 $S\left(T\right)=\frac{\mathsf{g_{up}}}{Q\left(T\right)}\,\frac{A}{8\pi c\hat{\nu}^{2}}\,e^{-hcE_{\mathrm{low}}/k_{B}T}\,\left(1-e^{-hc\hat{\nu}/k_{B}T}\right),$ (15)
@@
 with $T_{\mathrm{ref}}=296\,\mathrm{K}$ and $P_{\mathrm{ref}}=1\,\mathrm{bar}$.
-4.1.2 HITEMP
-According to HiTEMP, HITEMP provides $\hat{\nu}$, $S(T_{\mathrm{ref}})$ at $T_{\mathrm{ref}}=296\,\mathrm{K}$, and $E_{\mathrm{low}}$, from which
+IV.1.2 HITEMP
+According to Rothman et al. (2010), HITEMP provides $\hat{\nu}$, $S(T_{\mathrm{ref}})$ at $T_{\mathrm{ref}}=296\,\mathrm{K}$, and $E_{\mathrm{low}}$, from which
 $S\left(T\right)=S\left(T_{\mathrm{ref}}\right)\,\frac{Q\left(T_{\mathrm{ref}}\right)}{Q\left(T\right)}\,\frac{e^{-hcE_{\mathrm{low}}/k_{\mathrm{B}}T}}{e^{-hcE_{\mathrm{low}}/k_{\mathrm{B}}T_{\mathrm{ref}}}}\,\frac{1-e^{-hc\hat{\nu}/k_{\mathrm{B}}T}}{1-e^{-hc\hat{\nu}/k_{\mathrm{B}}T_{\mathrm{ref}}}}$ (17)
@@
 For some molecules, $\mathrm{H_{2}}$-broadened parameters $\gamma_{\mathrm{ref,,H_{2}}}$ and $\mathsf{n}_{\mathrm{H_{2}}}$ are available, allowing $\mathrm{H_{2}}$-background pressure broadening to be computed via Eq.
-(18) (tan2022).
+(18) (Tan et al., 2022).
 In what follows, we refer to air-broadened HITEMP as $\mathrm{HITEMP_{Air}}$ and $\mathrm{H_{2}}$-broadened HITEMP as $\mathrm{HITEMP_{H_{2}}}$.
 Nominally, air broadening parameters are not appropriate for brown dwarf atmospheres.
-However, the laboratory experiment by 2025ApJ...984...92H suggests that these parameters are not necessarily incorrect, so we include them in the comparison.
-4.1.3 Adopted Line Lists in This Study
-Given that brown dwarf atmospheres are dominated by $\mathrm{H_{2}}$, we adopt the ExoMol POKAZATEL for $\mathrm{H_{2}O}$ (POKAZATEL), ExoMol MM for $\mathrm{CH_{4}}$ (MM), and the ExoMol Coxon–Hajig line list for $\mathrm{HF}$ (Coxon-Hajig).
+However, the laboratory experiment by Hosokawa et al.
+(2025) suggests that these parameters are not necessarily incorrect, so we include them in the comparison.
+IV.1.3 Adopted Line Lists in This Study
+Given that brown dwarf atmospheres are dominated by $\mathrm{H_{2}}$, we adopt the ExoMol POKAZATEL for $\mathrm{H_{2}O}$ (Polyansky et al., 2018), ExoMol MM for $\mathrm{CH_{4}}$ (Yurchenko et al., 2024), and the ExoMol Coxon–Hajig line list for $\mathrm{HF}$ (Coxon and Hajigeorgiou, 2015).
 To assess the impact of the CO line list — which exhibits prominent absorption in the $K$ band — on the retrieval results, we consider the following three CO line list configurations:
 $\mathrm{ExoMol\,\,CO}$
-ExoMol CO line list Li2015 (Li2015), assuming an $\mathrm{H_{2}}$ background.
+ExoMol CO line list Li2015 (Li et al., 2015), assuming an $\mathrm{H_{2}}$ background.
 $\mathrm{HITEMP_{Air}\,\,CO}$
-HITEMP CO line list (HITEMPCO) with air-broadening parameters.
+HITEMP CO line list (Li et al., 2015) with air-broadening parameters.
 $\mathrm{HITEMP_{H_{2}}\,\,CO}$
-An $\mathrm{H_{2}}$-broadened version of HITEMP CO based on the semi-empirical models constructed on available and validated experimental measurements and theoretical predictions (tan2022).
+An $\mathrm{H_{2}}$-broadened version of HITEMP CO based on the semi-empirical models constructed on available and validated experimental measurements and theoretical predictions (Tan et al., 2022).
 For all three CO line lists, the same Voigt profile treatment is applied, so that the differences in the retrieval results originate only from the broadening parameters.
-The selection of molecular species was referenced from the $K$-band high-resolution spectral analysis of Luhman 16AB with IGRINS/Gemini South reported by Ishikawa2025.
+The selection of molecular species was referenced from the $K$-band high-resolution spectral analysis of Luhman 16AB with IGRINS/Gemini South reported by Ishikawa et al.
+(2025).
 We also searched for the isotopologue $^{13}\mathrm{CO}$, but it was not detected in the present analysis.
-This is consistent with Ishikawa2025, who did not detect $^{13}\mathrm{CO}$ in the same wavelength region.
+This is consistent with Ishikawa et al.
+(2025), who did not detect $^{13}\mathrm{CO}$ in the same wavelength region.
 Therefore, we did not include $^{13}\mathrm{CO}$ opacity in the spectral model.
-4.2 Results
+IV.2 Results
 Figure 3: Corner plots for the posterior samples obtained by fitting the spectra with models that use CO line lists from $\mathrm{ExoMol\,\,CO}$ (orange), $\mathrm{HITEMP_{Air}\,\,CO}$ (green), and $\mathrm{HITEMP_{H_{2}}\,\,CO}$ (blue).
@@
 Note that $\tau_{c}$ denotes the cumulative optical depth due only to cloud opacity.
-Blue, purple, and red lines are Sonora Diamondback (SD) TTPP models (Morley2024).
-Black curves in the right panels indicate condensation curves for the principal brown dwarf cloud species identified by Morley2024 (condensation curve formulas from Visscher2010; Wakeford2017).
+Blue, purple, and red lines are Sonora Diamondback (SD) TTPP models (Morley et al., 2024).
+Black curves in the right panels indicate condensation curves for the principal brown dwarf cloud species identified by Morley et al.
+(2024) (condensation curve formulas from Visscher et al.
+2010; Wakeford et al.
+2017).
 Table 2: Parameter estimates for Luhman 16A under each modeling configuration.
@@
 The retrieved TTPP profiles under the power-law + gray-cloud + $\mathrm{HITEMP_{H_{2}}\,\,CO}$ configuration are shown for each object in Figure 4.
-Across all line lists, the inferred temperature ranges in the regions of largest contribution are consistent with the Sonora diamondback models (Morley2024) at the literature effective temperatures.
+Across all line lists, the inferred temperature ranges in the regions of largest contribution are consistent with the Sonora diamondback models (Morley et al., 2024) at the literature effective temperatures.
 Here, $\tau_{c}$ denotes the cumulative optical depth due only to the gray-cloud opacity integrated downward from the top of the atmosphere, and the pressure level at $\tau_{c}=1$ corresponds to the cloud deck in our model.
@@
 $\mathrm{[(C+O)/H]}=\log_{10}\left(\frac{n_{\mathrm{C}}+n_{\mathrm{O}}}{n_{\mathrm{H}}}\right)-\log_{10}\left(\mathrm{C+O\over H}\right)_{\odot},$ (25)
-where the solar ratios $\left(\mathrm{X/H}\right)_{\odot}$ follow Asplund2021.
+where the solar ratios $\left(\mathrm{X/H}\right)_{\odot}$ follow Asplund et al.
+(2021).
 For both Luhman 16A and 16B, the VMR-inferred C/O is $\sim 0.67$ — slightly above solar (Figure 3).
@@
 Differences in the line shapes alter the inferred cloud-deck pressure and, in turn, the relative VMRs.
-To interpret the retrieved abundances, we computed thermochemical-equilibrium VMRs with FastChem including equilibrium condensation (fastchem; fastchem_cond).
-In these calculations, we adjusted the solar C, O, and F abundances of Asplund2021 by our retrieved $\mathrm{[X/H]}$ values (all other elements fixed to solar).
-Asplund2021 tabulate $A(\mathrm{X})\equiv\log_{10}(n_{\mathrm{X}}/n_{\mathrm{H}})+12$ with $A(\mathrm{H})=12$; accordingly we adopt
+To interpret the retrieved abundances, we computed thermochemical-equilibrium VMRs with FastChem including equilibrium condensation (Stock et al., 2018; Kitzmann et al., 2024).
+In these calculations, we adjusted the solar C, O, and F abundances of Asplund et al.
+(2021) by our retrieved $\mathrm{[X/H]}$ values (all other elements fixed to solar).
+Asplund et al.
+(2021) tabulate $A(\mathrm{X})\equiv\log_{10}(n_{\mathrm{X}}/n_{\mathrm{H}})+12$ with $A(\mathrm{H})=12$; accordingly we adopt
 $A(\mathrm{X})=A_{\odot}(\mathrm{X})+\mathrm{[X/H]},\quad\mathrm{X}\in\{\mathrm{C,O,F}\},$ (26)
 evaluated separately for Luhman 16A and 16B using their respective $\mathrm{[X/H]}$.
-Numerically, $A_{\odot}(\mathrm{C})=8.46$, $A_{\odot}(\mathrm{O})=8.69$, and $A_{\odot}(\mathrm{F})=4.40$ (Asplund2021).
+Numerically, $A_{\odot}(\mathrm{C})=8.46$, $A_{\odot}(\mathrm{O})=8.69$, and $A_{\odot}(\mathrm{F})=4.40$ (Asplund et al., 2021).
 The equilibrium VMRs computed with these inputs and the retrieved TTPP profiles (Figure 5) agree well for $\mathrm{CO}$, $\mathrm{H_{2}O}$, and $\mathrm{HF}$, indicating compatibility with chemical equilibrium under the inferred thermal structures.
-By contrast, the retrieved $\mathrm{CH_{4}}$ VMRs for both objects are about an order of magnitude lower than the FastChem predictions; we revisit this discrepancy in Section 5.3.
+By contrast, the retrieved $\mathrm{CH_{4}}$ VMRs for both objects are about an order of magnitude lower than the FastChem predictions; we revisit this discrepancy in Section V.3.
 Figure 5: Retrieved VMRs of $\mathrm{H_{2}O}$ (purple), $\mathrm{CO}$ (brown), $\mathrm{CH_{4}}$ (pink), and $\mathrm{HF}$ (yellow) for Luhman 16A (left) and Luhman 16B (right).
@@
 For comparison, thermochemical-equilibrium VMRs computed with FastChem, including equilibrium condensation, are shown as 5–95% credible bands.
-4.2.1 Comparison with Medium-Resolution Flux Density
+IV.2.1 Comparison with Medium-Resolution Flux Density
 Although the retrievals above were performed on normalized spectra, the forward models also provide absolute flux densities.
@@
 (Right) Luhman 16B.
-4.2.2 Impact of Photometric Variability in Luhman 16B
-Luhman 16A shows $2\%$ variability over a 7-hr baseline (Biller2024), and Luhman 16B exhibits $\sim$3% variability over 5 hr (Crossfield2014) and up to $\sim$6% over 7 hr (Biller2024).
-Such variability, common among L/T transition brown dwarfs including Luhman 16AB, is often attributed to heterogeneous silicate/iron clouds whose evolving rotational modulation produces temporal brightness changes (Radigan2014; Crossfield2014; Chen2024; Saumon_Marley2008).
-The CRIRES time series used here comprise 14 spectra ($\sim$5 hr) for each component (Crossfield2014) and show comparable normalized flux variability: 2.3% for A and 3.2% for B.
+IV.2.2 Impact of Photometric Variability in Luhman 16B
+Luhman 16A shows $2\%$ variability over a 7-hr baseline (Biller et al., 2024), and Luhman 16B exhibits $\sim$3% variability over 5 hr (Crossfield et al., 2014) and up to $\sim$6% over 7 hr (Biller et al., 2024).
+Such variability, common among L/T transition brown dwarfs including Luhman 16AB, is often attributed to heterogeneous silicate/iron clouds whose evolving rotational modulation produces temporal brightness changes (Radigan, 2014; Crossfield et al., 2014; Chen et al., 2024; Saumon and Marley, 2008).
+The CRIRES time series used here comprise 14 spectra ($\sim$5 hr) for each component (Crossfield et al., 2014) and show comparable normalized flux variability: 2.3% for A and 3.2% for B.
 Our baseline retrieval assumes a horizontally homogeneous (i.e., effectively averaged) atmosphere, neglecting cloud patchiness.
@@
 Hence, for the present data, the impact of ignoring cloud inhomogeneity appears subdominant to systematics associated with the line lists.
-This conclusion is consistent with the analysis of the $J$-band VLT/CRIRES+ spectrum by deRegt2025, who found that a patchy-cloud model is only weakly preferred for Luhman 16B and that the retrieval outcomes are largely insensitive to the inclusion of patchiness.
+This conclusion is consistent with the analysis of the $J$-band VLT/CRIRES+ spectrum by de Regt et al.
+(2025), who found that a patchy-cloud model is only weakly preferred for Luhman 16B and that the retrieval outcomes are largely insensitive to the inclusion of patchiness.
 Figure 7: C/O and cloud deck pressure for Luhman 16B from stacked and time-resolved spectra.
@@
 Error bars denote 90% credible region.
-In the left panel, the dotted horizontal line marks the solar C/O (=0.59=0.59; Asplund2021).
-5 Retrievals Using a Flexible Gaussian Process TTPP Profile
-Some high-resolution retrieval studies of brown-dwarf and exoplanet atmospheres adopt simple parameterizations of the TTPP profile, such as a power-law form (e.g., Kawahara2022; Kawashima2025; Kasagi2025).
+In the left panel, the dotted horizontal line marks the solar C/O (=0.59=0.59; Asplund et al., 2021).
+V Retrievals Using a Flexible Gaussian Process TTPP Profile
+Some high-resolution retrieval studies of brown-dwarf and exoplanet atmospheres adopt simple parameterizations of the TTPP profile, such as a power-law form (e.g., Kawahara et al., 2022; Kawashima et al., 2025; Kasagi et al., 2025).
 Although efficient, such profiles can be overly restrictive.
-More flexible parameterizations have therefore been proposed, including schemes that parameterize the temperature gradient in discrete pressure layers (Zhang2023) and $P$-splines (e.g., Line2015; deRegt2025).
+More flexible parameterizations have therefore been proposed, including schemes that parameterize the temperature gradient in discrete pressure layers (Zhang et al., 2023) and $P$-splines (e.g., Line et al., 2015; de Regt et al., 2025).
 However, these approaches still remain tied to specific functional forms and may fail to capture complex structures.
-As a potential future extension that takes advantage of the computational efficiency of HMC/NUTS in high-dimensional settings, [^1] The convergence time of conventional random Metropolis–Hastings scales as $D^{2}$ with $D$ being the number of parameters, whereas HMC scales as $D^{5/4}$ (HMC).
-There are also studies suggesting that NUTS scales as $D^{1/4}$ when initialized in the concentration region of the canonical Gaussian measure (Bou-Rabee2024).
-Kawahara2022 proposed a nonparametric TTPP profile model based on Gaussian processes (GP, 2023ARA&A..61..329A).
+As a potential future extension that takes advantage of the computational efficiency of HMC/NUTS in high-dimensional settings, [^1] The convergence time of conventional random Metropolis–Hastings scales as $D^{2}$ with $D$ being the number of parameters, whereas HMC scales as $D^{5/4}$ (Neal, 2011).
+There are also studies suggesting that NUTS scales as $D^{1/4}$ when initialized in the concentration region of the canonical Gaussian measure (Bou-Rabee and Oberdörster, 2024).
+Kawahara et al.
+(2022) proposed a nonparametric TTPP profile model based on Gaussian processes (GP, Aigrain and Foreman-Mackey, 2023).
 That work, however, did not explore its practical applicability in real retrievals, its sensitivity to other parameters, or its validation through injection–recovery tests.
 Here we refine the GP-based TTPP profile model into a more practical framework and demonstrate its performance through application to the retrieval of Luhman 16AB data.
-5.1 Formulation of the GP TTPP Profile
+V.1 Formulation of the GP TTPP Profile
 To flexibly represent temperature–pressure (TTPP) structures while maintaining numerical stability, we define the temperature profile as a GP in a whitened and logistically mapped latent space.
@@
 which we sample using NUTS.
-Relative to the GP formulation of Kawahara2022, our implementation introduces several changes aimed at practical convergence and physical regularization.
-First, while Kawahara2022 adopted a RBF kernel, we instead use a Matérn 3/2 kernel.
+Relative to the GP formulation of Kawahara et al.
+(2022), our implementation introduces several changes aimed at practical convergence and physical regularization.
+First, while Kawahara et al.
+(2022) adopted a RBF kernel, we instead use a Matérn 3/2 kernel.
 The Matérn kernel is less smooth than the RBF kernel, avoiding the overly smooth behavior inherent to the RBF assumption.
 We also infer the kernel hyperparameters (the amplitude $a_{u}$ and correlation length $\tau_{u}$) jointly with the atmospheric parameters, thereby removing the need to set them manually.
-In addition, instead of adopting a constant mean function as in Kawahara2022, we use a power-law form characterized by $(T_{0,\,m},\,\alpha_{m})$, whose parameters are inferred jointly with the GP.
-We find that a constant mean tends to force the retrieved profile to flatten toward the average value in the upper and lower layers, as is also hinted in figure 11 Kawahara2022.
+In addition, instead of adopting a constant mean function as in Kawahara et al.
+(2022), we use a power-law form characterized by $(T_{0,\,m},\,\alpha_{m})$, whose parameters are inferred jointly with the GP.
+We find that a constant mean tends to force the retrieved profile to flatten toward the average value in the upper and lower layers, as is also hinted in figure 11 Kawahara et al.
+(2022).
 The power-law mean model, on the other hand, better reflects the physically expected structure of self-luminous brown dwarfs while allowing the GP to capture deviations from it.
-5.2 Demonstration
+V.2 Demonstration
 Figure 8: Retrieval results for the simulation spectra, modeled with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a Gaussian process TTPP profile.
@@
 This behavior demonstrates that the GP TTPP approach can capture complex thermal structures where the data are informative, while appropriately inflating uncertainties where they are not.
-5.3 Application to Luhman 16AB
+V.3 Application to Luhman 16AB
 Figure 10: Retrieval results for the VLT/CRIRES high-resolution spectra of Luhman 16A (left) and Luhman 16B (right), modeled with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a Gaussian process TTPP profile.
@@
 (Right) Luhman 16B.
-Figure 13: Comparison between the retrieved volume mixing ratios and FastChem equilibrium predictions (solar abundances from Asplund2021). (Left) Luhman 16A. (Right) Luhman 16B.
+Figure 13: Comparison between the retrieved volume mixing ratios and FastChem equilibrium predictions (solar abundances from Asplund et al. 2021). (Left) Luhman 16A. (Right) Luhman 16B.
 We applied the GP TTPP retrieval to the CRIRES data of Luhman 16AB.
 Best-fit model spectra for both components are shown in Figure 10.
-Corner plots for the selected parameters are shown in Figure 11; the retrieved TTPP profiles are shown in Figure 12 along with the power-law profiles from Section 4; the retrieved parameter values are summarized in the rightmost columns of Tables 2 and 3.
+Corner plots for the selected parameters are shown in Figure 11; the retrieved TTPP profiles are shown in Figure 12 along with the power-law profiles from Section IV; the retrieved parameter values are summarized in the rightmost columns of Tables 2 and 3.
 Figure 12 shows that the broad trends of the inferred TTPP profiles are consistent with those obtained under the power-law parameterization.
@@
 In this regard, the GP TTPP profile provides a useful and conservative framework.
-6 Summary and Discussion
+VI Summary and Discussion
 Using ExoJAX, we performed high-resolution retrievals of Luhman 16AB under a gray-cloud assumption and inferred elemental abundance ratios.
 Taking the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP parameterization as a baseline, we then (i) repeated the analysis with alternative CO line lists and (ii) introduced a more flexible TTPP description based on GPs to assess the robustness of the inferred parameters.
-6.1 Comparison with Previous Work
+VI.1 Comparison with Previous Work
 In addition to the present work, several other high-resolution retrieval studies of Luhman 16AB have recently been reported.
@@
 Abundance
-The molecules included in our model, H2O, CH4, CO, and HF, have all been detected in the $H$- and $K$-band Gemini South/IGRINS high-resolution ($R\sim 28,000$) spectra of Luhman 16AB presented by Ishikawa2025.
+The molecules included in our model, H2O, CH4, CO, and HF, have all been detected in the $H$- and $K$-band Gemini South/IGRINS high-resolution ($R\sim 28,000$) spectra of Luhman 16AB presented by Ishikawa et al.
+(2025).
 Their study compared the spectra against forward models generated with PICASO, which incorporates radiative–convective equilibrium, Sonora Bobcat TTPP profiles, VIRGA cloud models, and nonequilibrium chemistry parameterized by a vertical eddy diffusion coefficient $K_{zz}$.
 The retrieved volume mixing ratios (VMRs) of the common molecules (their Figure 8) are broadly consistent with ours, but tend to be smaller, likely because they fixed the metallicity to solar whereas we inferred slightly super-solar values.
-Both H2O and HF have also been detected in the CRIRES+/VLT $J$-band high-resolution retrieval by deRegt2025.
-Their modeling adopted a TTPP parameterization based on the temperature gradient at five knots (as in Zhang2023), and clouds were treated as gray opacity decreasing with a power-law of pressure from the cloud base upward.
+Both H2O and HF have also been detected in the CRIRES+/VLT $J$-band high-resolution retrieval by de Regt et al.
+(2025).
+Their modeling adopted a TTPP parameterization based on the temperature gradient at five knots (as in Zhang et al., 2023), and clouds were treated as gray opacity decreasing with a power-law of pressure from the cloud base upward.
 Despite these differences, the retrieved molecular abundances are consistent with ours.
-Neither study provides explicit estimates of the C/O ratio: Ishikawa2025 fixed the metallicity and C/O to solar, and in the spectral range analyzed by deRegt2025, no carbon-bearing molecules were detected.
+Neither study provides explicit estimates of the C/O ratio: Ishikawa et al.
+(2025) fixed the metallicity and C/O to solar, and in the spectral range analyzed by de Regt et al.
+(2025), no carbon-bearing molecules were detected.
 Cloud
-Ishikawa2025 also detected H2 lines.
+Ishikawa et al.
+(2025) also detected H2 lines.
 By comparing the inferred H2 column density with the interstellar gas-to-dust ratio, they suggested that dust clouds may reside deeper than the H2 absorption layer, located at $\sim 3\,\mathrm{bar}$ (or $\log_{10}P\sim 0.5$) in both components A and B.
 Their self-consistent models likewise imply that clouds are present at similar or deeper pressures relative to the H2 line-forming region.
-In the power-law cloud model of deRegt2025, the cloud opacity is dominated by contributions around $\log_{10}P\sim 0.5$–1, and this estimate is robust to assumptions about patchiness.
+In the power-law cloud model of de Regt et al.
+(2025), the cloud opacity is dominated by contributions around $\log_{10}P\sim 0.5$–1, and this estimate is robust to assumptions about patchiness.
 In our analysis as well, the inferred cloud top pressures consistently fall within $\log_{10}P\sim 0.5$–1, broadly agreeing with these results.
 On the other hand, the quantitative estimates depend on assumptions about the TTPP profile and line lists, which in turn propagate into the inferred C/O ratio as discussed in the next subsection.
-6.2 $\mathrm{C/O}$ Ratios
+VI.2 $\mathrm{C/O}$ Ratios
 Figure 14 compares $\mathrm{C/O}$ inferred with each CO line list configuration and with the GP TTPP model (using $\mathrm{HITEMP_{H_{2}}\,\,CO}$) to literature values for other brown dwarfs and super Jupiters.
@@
 By contrast, the assumed TTPP profile (power-law versus GP) does not yield a significant difference in $\mathrm{C/O}$.
-An epoch-by-epoch analysis of the Luhman 16B time series likewise reveals no statistically significant impact from cloud-cover variability on the retrieved $\mathrm{C/O}$ or molecular VMRs (Section 4.2.2).
+An epoch-by-epoch analysis of the Luhman 16B time series likewise reveals no statistically significant impact from cloud-cover variability on the retrieved $\mathrm{C/O}$ or molecular VMRs (Section IV.2.2).
 We therefore conclude that, in our framework, the principal systematic on $\mathrm{C/O}$ arises from the CO line list, whereas the TTPP shape and photometric variability are not dominant.
@@
 While differences in formation pathways (i.e., gravitational instability vs.
-core accretion) remain a possible explanation, the contrast could also reflect differences in the efficiency of oxygen sequestration by silicate (and other) clouds, as emphasized in previous low-resolution studies (e.g., Line2015; Line2017): more efficient rainout in field brown dwarfs would raise the gas-phase C/O, whereas weaker sequestration in super Jupiters would leave C/O closer to the stellar value.
+core accretion) remain a possible explanation, the contrast could also reflect differences in the efficiency of oxygen sequestration by silicate (and other) clouds, as emphasized in previous low-resolution studies (e.g., Line et al., 2015, 2017): more efficient rainout in field brown dwarfs would raise the gas-phase C/O, whereas weaker sequestration in super Jupiters would leave C/O closer to the stellar value.
 A decisive test will require a larger, uniformly analyzed sample of both planets and brown dwarfs, so that inter-study systematics are minimized and C/O can be compared on equal footing.
 Building on such a uniform baseline, joint analyses of high-resolution spectra (to constrain TTPP) and JWST coverage (to probe silicate or other cloud composition and vertical structure) can further mitigate cloud–TTPP–chemistry degeneracies, thereby providing a more direct test of the efficiency of oxygen sequestration by rainout.
-We return to the prospects and implementation of such joint retrievals in Section 6.4.
+We return to the prospects and implementation of such joint retrievals in Section VI.4.
 Figure 14: Comparison of $\mathrm{C/O}$ for Luhman 16AB derived in this work with $\mathrm{C/O}$ values reported for other brown dwarfs and directly-imaged super Jupiters from high-resolution spectroscopic analyses (Table 4).
 Abbreviations: PLTP = power-law TTPP profile; GPTP = Gaussian-process TTPP profile.
-6.3 Usefulness and Caveats of the GP TTPP Model
+VI.3 Usefulness and Caveats of the GP TTPP Model
 For both simulations and real data, the GP TTPP model captures structures that are more complex than a simple power-law and yields a more faithful assessment of uncertainties.
 In the present application, these differences have only a modest impact on bulk parameters such as $\mathrm{C/O}$.
-However, as discussed in Section 5.3, comparisons between retrieved molecular abundances and chemical-equilibrium predictions are sensitive to how accurately TTPP uncertainties are estimated; in this context, the GP framework is important because it enables predictive uncertainties to faithfully reflect the spectra’s information content, remaining conservative where the data are weakly informative.
+However, as discussed in Section V.3, comparisons between retrieved molecular abundances and chemical-equilibrium predictions are sensitive to how accurately TTPP uncertainties are estimated; in this context, the GP framework is important because it enables predictive uncertainties to faithfully reflect the spectra’s information content, remaining conservative where the data are weakly informative.
 The GP approach is also sensitive to residual model–data mismatches and thus shows promise for diagnosing deficiencies in current models, potentially pointing to additional physics.
-On the other hand, as indicated in Section 5.3, degeneracies between the TTPP shape and cloud opacity can also allow nonphysical temperature structures if left unconstrained.
+On the other hand, as indicated in Section V.3, degeneracies between the TTPP shape and cloud opacity can also allow nonphysical temperature structures if left unconstrained.
 Mitigations include imposing weak physical regularization (e.g., monotonicity or stability constraints) on flexible TTPP profiles and/or extending the analysis to broader wavelength coverage where cloud signatures are stronger.
-The latter strategy can help balance flexibility with physical plausibility; see also Section 6.4.
+The latter strategy can help balance flexibility with physical plausibility; see also Section VI.4.
 Figure 15: Comparison between the medium-resolution spectra predicted by extrapolating our high-resolution retrievals and the JWST observations.
 Shaded bands show the 90% credible intervals of the predicted spectra computed with the $\mathrm{HITEMP_{H_{2}}\,\,CO}$ line list and a power-law TTPP profile (blue: Luhman 16A; orange: Luhman 16B).
-Points show the corresponding JWST medium-resolution spectra from Biller2024; Chen2025 (blue: A; orange: B).
-6.4 Cloud Degeneracies and Joint Analysis with JWST
+Points show the corresponding JWST medium-resolution spectra from Biller et al.
+(2024); Chen et al.
+(2025) (blue: A; orange: B).
+VI.4 Cloud Degeneracies and Joint Analysis with JWST
 Throughout this work, we find that the inferred cloud-top pressure is sensitive to assumptions about the TTPP profile and the adopted line list.
@@
 HD 33632 Ab FBD Y $0.58\pm 0.14$ $46\pm 8$ $1628^{+109}_{-97}$ (4), (5)
-Gl 229 B FBD Y $0.77\pm 0.01$ $71.47^{+0.57}_{-0.61}$ccGl 229 B has been reported to be a binary, based on high–angular-resolution observations with GRAVITY/VLTI together with radial-velocity monitoring with CRIRES+/VLT (Xuan2024b).
+Gl 229 B FBD Y $0.77\pm 0.01$ $71.47^{+0.57}_{-0.61}$ccGl 229 B has been reported to be a binary, based on high–angular-resolution observations with GRAVITY/VLTI together with radial-velocity monitoring with CRIRES+/VLT (Xuan et al., 2024b).
 $850\pm 50$ (6), (7)
@@
 Note.
-— References—(1) deRegt2024; (2) Xuan2022; (3) Xuan2024c; (4) Hsu2024; (5) Currie2020; (6) Kawashima2025; (7) Calamari2022; (8) Wang2022; (9) Mulder2025; (10) Schlawin2017; (11) Zapatero2014; (12) Filippazzo2015; (13) Costes2024; (14) Franson2022; (15) Johnson-Groh2017; (16) Picos2024; (17) Scholz2005; (18) Cooper2024; (19) Morales2021; (20) Zhang2021; (21) Vos2022; (22) Zhang2024b; (23) Currie2023; (24) Landman2024; (25) Lacour2021; (26) Chilcote2017; (27) Picos2025; (28) Schwarz2016; (29) Zhang2024a; (30) Wood2023; (31) Xuan2024a; (32) Gandhi2025; (33) Bonnefoy2010
+— References—(1) de Regt et al.
+(2024); (2) Xuan et al.
+(2022); (3) Xuan et al.
+(2024c); (4) Hsu et al.
+(2024); (5) Currie et al.
+(2020); (6) Kawashima et al.
+(2025); (7) Calamari et al.
+(2022); (8) Wang et al.
+(2022); (9) Mulder et al.
+(2025); (10) Schlawin et al.
+(2017); (11) Zapatero Osorio et al.
+(2014); (12) Filippazzo et al.
+(2015); (13) Costes et al.
+(2024); (14) Franson et al.
+(2022); (15) Johnson-Groh et al.
+(2017); (16) González Picos et al.
+(2024); (17) Scholz et al.
+(2005); (18) Cooper et al.
+(2024); (19) Morales-Gutiérrez et al.
+(2021); (20) Zhang et al.
+(2021); (21) Vos et al.
+(2022); (22) Zhang et al.
+(2024b); (23) Currie et al.
+(2023); (24) Landman et al.
+(2024); (25) Lacour et al.
+(2021); (26) Chilcote et al.
+(2017); (27) González Picos et al.
+(2025); (28) Schwarz et al.
+(2016); (29) Zhang et al.
+(2024a); (30) Wood et al.
+(2023); (31) Xuan et al.
+(2024a); (32) Gandhi et al.
+(2025); (33) Bonnefoy et al.
+(2010)